### v2

# Data Trade and Consumer Privacy

Abstract
This paper studies optimal mechanisms for collecting and trading data. Consumers benefit from revealing information about their tastes to a service provider because this improves the service. However, the information is also valuable to a third party as it may extract more revenue from the consumer in another market called the product market. The paper characterizes the constrained optimal mechanism for the service provider subject to incentive feasibility. It is shown that the service provider sometimes sells no information or only partial information in order to preserve profits in the service market. In a general setup, the service provision distortion and no-price discrimination in the product market are exclusive. Moreover, a ban on data trade may reduce social welfare because it makes it harder to price discriminate in the product market.
Keywords: Information market, Privacy, Price discrimination, Mechanism design, Information design
JEL Codes: D82, D83
1 Introduction
The digital economy is deeply rooted in the sharing and use of personal data. Meanwhile, the transmission of personal data across markets brings concerns on privacy, [^1] For instance, personal data of Facebook users was collected by British consulting firm Cambridge Analytica in the 2010s. The privacy issues attract attention from both the public and regulatory authorities. The General Data Protection Regulation was put into effect in Europe in 2018, the California Consumer Privacy Act went into effect in 2020 and China’s Personal Information Protection Law was in force in late 2021. which can bite: The privacy concerns induced by data transmission, such as data trade, can adversely affect the data generating activities (e.g., data sharing, service usage) in related markets. The special feature of data endows it with a cross-market role.
Data is usually collected in one market and is also valuable in other markets. For example, Amazon Music, a music service provider, may collect taste data when users are enjoying the service of music recommendations. Amazon Music can trade the user data to a headphone seller (firm, producer, or data buyer) and the latter can use the data to price against users’ interests. As consumers have become increasingly aware of how their data can be collected, traded, and used, their willingness to use Amazon Music service and to share their data will be affected. With the cross-market role of data trade, the intermediary (service provider or platform) would design mechanisms for both collecting and trading data. What are the optimal mechanisms? How do the data trade and its cross-market effect shape the equilibrium privacy, surplus allocation, and social welfare?
To address these questions, this paper develops an integrated model including three sorts of interactions: (1) the intermediary’s data collection from consumers by providing service in the service market; (2) data trade between the intermediary and a firm in the data market; (3) pricing game between the firm and consumers in the product market. Figure 1 illustrates an overview of these interactions. [^2] The interactions in Figure 1 fits into many market scenarios such as Borrower-Fintech-Lender. In the example of Amazon Music, consumers share their taste information to obtain better music recommendations in the service market. Amazon Music sells the data to a headphone seller, who will sell headphones to consumers in the headphone market. We use the framework to study the intermediary optimal mechanisms of data collection and trade.
Our framework enables formal investigations on data sourcing and data trade in a comprehensive way. This contributes to the recent advocacy of integrated models to comprehensively understand the information markets (Bergemann and Bonatti, 2019) with the seminal envision back to Marschak (1968). Many works have separately investigated some specific elements, like data broker’s information selling mechanism or privacy protection, and less has offered a comprehensive investigation on data collection, data trade, and data usage. The holistic study of collection, trade, and usage of data and their effects and regulation implications cannot be easily discussed in divided models.
intermediary consumers firm data tradepaymentdata collectionpaymentservicepaymentproduct
Figure 1: The overview of the modeling ingredients and interactions among players.
In the framework, consumers who have private information about horizontal characteristics (tastes) benefit from sharing information with the intermediary by receiving a better-matched service but at the risk of the shared information being sold to a third party, the producer who can price against consumers using the bought data. In detail, consumers’ tastes are $H$ or $L$ (e.g., classical or rock music fans). Each consumer’s utility is from consuming service and product. The service provided to consumers may generate a mismatch as the intermediary does not know consumers’ true tastes. Each consumer also has unit demand for the product. The producer, a monopolist in the product market, does not know consumers’ willingness to pay for the product, which is related to consumers’ tastes. Hence, the taste data collected by the intermediary is valuable for the producer to set product prices. The intermediary can sell consumer data to the producer and it is the unique data source. The intermediary specifies (and commits to) a contract that consists of service provision, service fee, privacy policy (data trade), and data fee to consumers and the producer. Consumers and the producer decide whether to accept. Understanding that the data trade may discourage the service usage and data sharing of consumers, the intermediary’s contract would trade off between benefits of trading data and costs of hindering consumers’ data sharing and service usage.
We characterize the intermediary constrained optimal mechanisms. We first consider, given an intermediary’s contract, the decisions of the producer and consumers which impose participation and incentive constraints on the intermediary’s design problem. First, the producer buys data for a binary pricing decision, the binary-signal data works as product price recommendations subject to the obedience constraints (Bergemann and Morris, 2016). That is, the producer follows price recommendations of the traded data. Thus, data trade affects the product price and hence consumers surplus. The intermediary’s data offer takes a leave-it-or-take-it form. The payment of data trade makes the producer willing to accept the intermediary’s data offer. Second, we turn to consumers’ decisions. As the privacy policy affects consumers surplus in the product market, consumers’ incentive to reveal their types in the service market is then affected by the privacy policy. The choices of privacy policy and service make consumers willing to accept the offer and also to reveal their types.
One key feature of the intermediary optimal mechanisms is that the optimal informativeness of data trade depends on comparisons of the horizontal differentiation level in the data-sourcing market (captured by the variation in service tastes) and vertical differentiation level in data-using market (captured by the variation in willingness to pay for product). The optimal data trade sells no or partial information when the horizontal differentiation level is low relative to the vertical one. In this case, as the service mismatch is small, consumers have less incentive to share their data to get the right service. The intermediary commits to selling less data to maintain consumers’ incentives to reveal information. Otherwise, suppose that the intermediary sells full information to the producer. Although the intermediary obtains most consumers surplus in the product market, it has to incentivize consumers’ information revelation by reducing service fees and hence loses the service profit. When the horizontal differentiation level is relatively low, the service profit losses outweigh the extra surplus gain in the product market. Hence, the intermediary would prefer to sell less data to preserve the service profits. In contrast, when the horizontal differentiation level is relatively high, the intermediary can sell full information. The service profit losses due to incentive distortions are small even if the intermediary sells full information. Under the full information data trade, consumers are price discriminated against in the product market.
The intermediary’s tradeoff holds in both the binary and general service space, but with different thresholds for the comparisons. As the service provision affects consumers’ incentives, we investigate a setup of general service in Section 5, in which the intermediary has more flexible service provisions and the service provided to the type-L consumers is downward distorted under some conditions. [^3] Section 4 assumes a binary service space to focus on the role of data trade, in which there is no service distortion to consumers in the intermediary optimal mechanisms. However, with these conditions, no-price discrimination in the product market cannot appear because the no-information data trade is not part of the optimal mechanisms. In other words, no-discrimination in product prices and distortion in the service provision are exclusive. In addition, in the optimal mechanisms, the intermediary extracts all the consumer surplus which is endogenous to data trade. This is because the data trade allows the intermediary to take back the consumer surplus of the product.
Using the framework, we evaluate the effects of regulations on the data market. Our results suggest that the data trade ban does not improve social welfare. This happens because without any data it is difficult for the producer to price discriminate consumers. Suppose the producer believes that the product market consists of more type-H consumers, then without extra data the producer would set a high price and only sell to the type-Hs. Under the optimal data trade, the producer can extract the type-L consumers surplus by price discrimination, who are otherwise excluded from the market.
The findings have several implications. First, the data market itself can protect consumer privacy, for example, when the horizontal differentiation level in the data-sourcing market is low. Second, regulators need a holistic viewpoint over regulations on the data market. If regulators focus on only privacy protection or data trade, policies that aim at improving welfare may decrease it. Third, public monitoring of the intermediary’s implementation of data policy is needed. Otherwise, there can be a commitment issue with the intermediary’s privacy policy. The no-information data trade is optimal under some conditions from the cross-market perspective. But from an inter-temporal viewpoint, in practice, after the consumers reveal their data to the intermediary, it has incentives to trade consumers’ information for profits. Fourth, the co-existence of consumer data sharing and informative data trade in the optimal mechanisms rationalizes the so-called digital “privacy paradox”. [^4] It refers to the general discrepancies between consumers’ active data sharing behaviors and their self-stated privacy concerns. Several surveys, experimental, and empirical works are about this observation, see Gross and Acquisti (2005); Goldfarb and Tucker (2012); Athey et al. (2017); Chen et al. (2021) for example. The consumers’ information sharing and intermediary’s data selling as an equilibrium outcome are consistent with the recent evidence. Ichihashi (2023) attributes the paradox to consumers’ decreasing marginal privacy cost in a dynamic interaction with the platform.
Related Literature.
The paper is related to several strands of literature: information market, the economics of privacy, price discrimination, information design, and mechanism design.
First, our paper adds to the literature on information markets (see recent surveys in Bergemann and Bonatti (2019) and Bergemann and Ottaviani (2021)). Admati and Pfleiderer (1990) focus on a monopolist that sells information about an asset and Bergemann and Bonatti (2015) study a data broker who sells cookies for targeting markets. Babaioff et al. (2012) and Bergemann et al. (2018) study optimal mechanisms for selling information in different contracting environments. Ichihashi (2021) and Bonatti et al. (2024) respectively investigate the competition among data intermediaries and data buyers. The majority of literature treats data sourcing as given. Few of these works investigate the mechanism designs for both collecting and selling information, which is the main attempt of the present paper. This paper highlights the cross-market effect of data trade.555 Consumer behaviors, especially as main data sources, can be distorted by the data trade. Argenziano and Bonatti (2023b) analyze strategic behaviors of privacy-conscious consumers. Several works show that the consumers’ manipulation incentives reduce the amount of information transmission in equilibrium, e.g., Frankel and Kartik (2022), Bonatti and Cisternas (2020).
Several works share a similar feature of tradeoffs between data exploitation and user privacy. Fainmesser et al. (2023) study the tradeoff between data collection and data protection but without a role of intermediary’s data trade, and hence the privacy is not endogenous to the data trade. [^6] They assume that the personal information (consumer activity such click) is verifiable. The privacy level is the protection investment. Argenziano and Bonatti (2023a) micro founds the privacy preferences. In our paper consumer privacy is based on the equilibrium data sale. Jullien et al. (2020) consider the tradeoff between data exploitation and user activity. Dosis and Sand-Zantman (2023) focus on a tradeoff between data monetization and processing. In these works, consumer information is verifiable. Our paper adopts the mechanism design approach because consumers control the information and information is unverifiable, which is more about consumers’ private characteristics than their verifiable activities, like browsing data. In addition, in our paper the tradeoff comes from the cross-market effect of data, relying on comparisons between the differentiation levels in data-sourcing and data-using markets.
Related to the cross-market effect of data, Doval and Skreta (Forthcoming) study the inter-temporal role of data with a focus on the tradeoff between product personalization and price discrimination in a dynamic model of upstream downstream. Rather, our paper investigates the cross-market role of data in a static environment. They find the no-price discrimination result under equal weight on profits of different periods in the commitment environment. No-discrimination can show up in the product market only if there is no distortion in service provisions to consumers in the data sourcing market.
This paper also belongs to the literature on price discrimination. Bergemann et al. (2015) study how market segmentation can support an arbitrary surplus distribution between consumers and a monopolist. The endogenous market segment also arises from consumers’ equilibrium disclosure about their characteristics in Ichihashi (2020) and Ali et al. (2023) and data broker’s creation and sale of market segments for profits in Yang (2022). This paper complements the literature by focusing on the critical role of the intermediary in both collecting data and trading data.
The price discrimination endogenous to the information is related to the economics of privacy with seminal works of Taylor (2004) and Calzolari and Pavan (2006); see Acquisti et al. (2016) and Goldfarb and Que (2023) for recent surveys. The privacy concerns arise from sellers’ learning about consumers’ type (Villas-Boas, 2004; Conitzer et al., 2012) and buyers’ information revelation to sellers (Hidir and Vellodi, 2021; Ichihashi, 2020). Argenziano and Bonatti (2023a) study the data linkages instead of the intermediary’s data selling. Ali et al. (2023) combine the consumer information disclosure and personalized pricing. In dynamic environments, the privacy concerns on discrimination induce a ratchet effect (Laffont and Tirole, 1988). The ratchet effect can bite all the information value as the data collecting intermediary gets better off committing to offering full privacy (Calzolari and Pavan, 2006). Doval and Skreta (Forthcoming) find in a limited commitment environment that the price discrimination induces product line gap due to this effect. Rhodes and Zhou (2021) study consumer privacy choice in the context of personalized pricing. Our paper extends the literature by endogenizing the privacy policy as a part of the intermediary’s holistic design of information collection and sale.
In addition, it is related to the broad literature of information design (Kamenica and Gentzkow, 2011; Bergemann and Morris, 2019) and mechanism design. [^7] See Mussa and Rosen (1978), Myerson (1981) for classic works and Bergemann and Välimäki (2006) for a survey on information in mechanism design. In our paper consumers, as the data sources, have private unverifiable information. Hence, we rely on the mechanism design approach to incentivize consumers’ information revelation and the privacy policy or data trade is one of the intermediary’s design elements. Bergemann et al. (2018) study the mechanism of selling information. Bonatti et al. (2024) investigate data selling to competing buyers and Doval and Skreta (Forthcoming) study personalization of product and prices also in a mechanism design framework. We embed the information design approach to selling information into the intermediary’s mechanism design problem of incentivizing consumers’ information revelation to highlight the cross-market role of data trade. [^8] Yoder (2022) explores the contract to incentivize agent’s information acquisition and Georgiadis and Szentes (2020) study the designs of wage and performance evaluation policies as incentive tools. In our paper, one feature of optimal mechanisms is that the intermediary, as the designer, can extract all consumers surplus as the consumers surplus is endogenous.
Lastly, although this paper abstracts from data externalities, the finding that data trade improves social welfare is related to the works studying externality in the data market. Acemoglu et al. (2022) find that banning the data market improves welfare as the over-sharing of data in the market induced by data externality disappears. The welfare loss comes from the firm’s over-collection of consumer data in Bergemann et al. (2022) and Choi et al. (2019).
Outline.
The rest of this paper is organized as follows. Section 2 sets up the model. Section 3 takes some preliminary analyses of individual decisions. Section 4 characterizes the intermediary optimal mechanism under binary service and discusses regulations of the data market. Section 5 extends to a general service space. Section 6 presents some discussions and section 7 concludes. Omitted proofs and additional results are relegated to Appendices.
2 Model
There are three types of players: one intermediary (it), one producer (he), and one unit mass of consumers (she). Consumers, the intermediary, and the producer interact in three markets: a service market, a product market, and a data market. In the service market, the intermediary collects data about consumers’ type by providing consumers service. The data can be traded in the data market from the intermediary to the producer, who sells products to consumers in the product market.
Each consumer has a type $\theta\in\left\{L,H\right\}$ with $L,H\in\mathbb{R}$. The type is a horizontal taste parameter. [^9] In this sense notations $H$ and $L$ have no essential meanings. One can relabel them as “A”“B”, e.t.c. Consumers’ type is realized from a prior distribution $\mu_{0}=\text{Pr}\left(\theta=H\right)\in[0,1]$, which is common knowledge. Each consumer knows her type, but neither the producer nor the intermediary observes consumers’ true types.
Consumers’ utility relies on the service and the product. A consumer of type $\theta$ who consumes both the service and product will get
$V-\underbrace{\left(x-\theta\right)^{2}}_{\text{ mismatch in service}}-f+\underbrace{\max\left\{v_{\theta}-p,0\right\}}_{\text{product market}}$
which consists of two parts. First, the recommended service by the intermediary generates a baseline utility level $V\in\mathbb{R}$, but there can be recommendation bias or mismatch between a recommendation $x$ and consumer true taste $\theta$, which is measured by the squared item. [^10] The squared form specifies only the mismatch formula in the equilibrium. Consumers need to pay the service fee $f$ to get service. [^11] The fee $f$ can be negative in the setup, which means that the intermediary pays the consumers. As we will see later the optimal $f$ will be always positive if $V$ is sufficiently large.
Second, each consumer has unit demand for the product and decides whether to purchase the product or not. Consumers of $\theta$ purchase the product if and only if $v_{\theta}-p\geq 0$ where $p$ denotes the product price and $v_{\theta}\in\mathbb{R}$ is the willingness to pay for the product of type $\theta$ consumers. We assume $v_{L}\neq v_{H}$ so that the horizontal type is informative about the willingness to pay. [^12] In contrast, if $v_{L}=v_{H}$ then data on consumers’ type is not valuable to the producer. It is without loss to assume $v_{L}<v_{H}$.
The producer is a monopolist in the product market. As the product demand relies on the producer’s information about consumers’ willingness to pay, the product revenue is
$p\times\text{Pr}\left(v\geq p|\mathcal{I}\right)$
where $\text{Pr}\left(v\geq p|\mathcal{I}\right)$ is the total mass of consumers who purchase the product at a price $p$ conditional on the producer’s information about consumer willingness to pay, $\mathcal{I}$.
The intermediary plays in both service and data markets by offering a contract of data collection and data trade. In the service market the intermediary is a monopolist who provides consumers service to collect consumers data on willingness to pay. The intermediary sells data via a data offer and the producer buys data for product pricing decision. The intermediary is the unique data source. Its profit comes from service provision and data sale.
By the revelation principle (Myerson, 1979), the intermediary designs a direct mechanism which specifies: (1) a service menu for consumers including the service provision $x:\{L,H\}\to\mathcal{X}$ where $\mathcal{X}$ is the set of service; and the service fee $f:\{L,H\}\to\mathbb{R}$; and (2) a data offer for the producer consisting of the disclosure policy captured by a pair of distributions over disclosed signals about consumers types $\pi:\left\{L,H\right\}\to\Delta\mathcal{S}$ where $\mathcal{S}$ is the set of possible signals; and the lump-sum data fee $T\in\mathbb{R}$. Consumers observe all the specifications. If a consumer reports $\theta^{\prime}$, then a signal will realize according to $\pi\left(s|\theta^{\prime}\right)$ in the product market. In sum, the intermediary offers a contract of data collection and data trade of which $\left(\mathcal{X},x,f\right)$ and ($\left(\left\{\mathcal{S},\pi\right\},T\right)$ specify the part of service market and data market respectively.
Timing.
Consumers privately observe type $\theta$. The intermediary posts the entire contract $\{\left(\mathcal{X},x,f\right),\left(\left\{\mathcal{S},\pi\right\},T\right)\}$ including the service menu in the service market and data offer in the data market. Then consumers and the producer make acceptance decisions. If consumers accept, they report type $\theta^{\prime}$ and then get service $x_{\theta^{\prime}}$ and pay $f_{\theta^{\prime}}$. If consumers reject, they get the outside option in the service market. If the producer accepts, after data trade, in the product market a signal $s$ about consumer’s type is realized according to $\pi$. The producer observes $s$ and then sets the product price. Consumers decide whether to purchase the product or not. If the producer rejects the data offer, then the product pricing is made under the prior information $\mu_{0}$. The timing of events is summarized in Figure 2.
consumers observe typeintermediary posts entire contract of service and dataconsumers and producer decide whether to acceptproducer sets product priceconsumers decide whether to purchase
Figure 2: Timing of the events.
Discussion about Modeling Assumption.
A key modeling assumption is the perfect correlation between consumers’ taste and their willingness to pay. That is to say, $v_{L}\neq v_{H}$ in the setup. Otherwise, the data about consumers’ tastes is not valuable to the producer. The relation between consumers’ taste parameter and their willingness to pay can be represented by $\beta_{\theta}=\text{Pr}\left(v_{H}|\theta\right)=1-\text{Pr}\left(v_{L}|\theta\right)$. The assumption reads $(\beta_{H}=1,\beta_{L}=0)$ or $(\beta_{H}=0,\beta_{L}=1)$. If we relax the assumption by $\beta_{H},\beta_{L}\in\left(0,1\right)$ with $\beta_{H}\neq\beta_{L}$, the main results don’t change.
3 Preliminaries
The intermediary’s total payoff consists of profit in the service market $\mu_{0}f_{H}+\left(1-\mu_{0}\right)f_{L}$ and profit in the data market $T$. That is,
$\mu_{0}f_{H}+\left(1-\mu_{0}\right)f_{L}+T.$ (1)
Before digging into the intermediary optimal mechanism, we first consider the decisions of consumers and the producer after observing the intermediary’s contract for both service and data markets and before observing a signal realization.
Interim Payoff of Producer.
In the product market, a consumer $\theta$ will purchase the product if and only if $v_{\theta}\geq p$ at the product price $p$. After observing a signal $s\in\mathcal{S}$ from the traded data $\pi$, the producer’s posterior belief is $\mu_{s}\equiv\text{Pr}(\theta=H|s)$ or $\text{Pr}(v=v_{H}|s)$
$\mu_{s}=\frac{\mu_{0}\pi\left(s|H\right)}{\mu_{0}\pi\left(s|H\right)+\left(1-\mu_{0}\right)\pi\left(s|L\right)}$ (2)
Hence, for each $s\in\mathcal{S}$ the producer’s demand function reads
$D\left(p;s\right)\equiv\text{Pr}\left(v\geq p|s\right)=\left\{\begin{array}[]{cc}0&\text{if }p>v_{H}\\ \mu_{s}&\text{if }v_{L}<p\leq v_{H}\\ 1&\text{if }p\leq v_{L}\end{array}\right.$
So the producer’s product pricing decision after observing $s$ is
$p_{s}\in\left\{\begin{array}[]{cc}\{v_{H}\}&\text{if }\mu_{s}>\frac{v_{L}}{v_{H}}\\ \{v_{L},v_{H}\}&\text{if }\mu_{s}=\frac{v_{L}}{v_{H}}\\ \{v_{L}\}&\text{if }\mu_{s}<\frac{v_{L}}{v_{H}}\end{array}\right.$ (3)
and the corresponding product revenue is $\max\{v_{H}\mu_{s},v_{L}\}$.
At the time of data trade, the producer’s expected product revenue from accepting the data trade offer ($(\{\mathcal{S},\pi\},T)$ is
$R\left(\pi\right)=\mathbb{E}_{\pi}\left[\max\left\{v_{H}\mu_{s},v_{L}\right\}\right]$ (4)
The producer’s payoff is $R\left(\pi\right)-T$ if accepting the data offer.
Interim Payoff of Consumers.
Given the intermediary’s contract, consumers of type $\theta$’s total interim payoff from accepting the offer and reporting $\theta^{\prime}$ is
$u\left(\theta,\theta^{\prime};\pi\right)=V-\left(x_{\theta^{\prime}}-\theta\right)^{2}-f_{\theta^{\prime}}+\mathbb{E}_{\pi\left(s|\theta^{\prime}\right)}\left[\max\left\{v_{\theta}-p_{s},0\right\}\right]$ (5)
where if a consumer of $\theta$ reports $\theta^{\prime}$, the recommendation bias is measured by $\left(x_{\theta^{\prime}}-\theta\right)^{2}$, and the continuation value of accepting is $\mathbb{E}_{\pi\left(s|\theta^{\prime}\right)}\left[\max\left\{v_{\theta}-p_{s},0\right\}\right]$.
The following investigates, given an intermediary’s contract, the optimal decisions of the producer and consumers, which then impose constraints on the intermediary’s design problem.
3.1 Producer Optimality
The producer buys data to make the product pricing decision. As there are two product prices $\left\{v_{L},v_{H}\right\}$, it is without loss of generality that the data consists of two signals $\mathcal{S}=\left\{l,h\right\}$. The data trade takes the form of price recommendation for the producer (Bergemann and Morris, 2016). Let the producer charge $v_{L}$ after observing signal $l$ and $v_{H}$ after observing signal $h$. That is,
$p_{l}=v_{L}\text{, }p_{h}=v_{H}$
Given a data trade $\pi$, the producer charges a consumer of type $\theta$ a product price $v_{L}$ with probability $\pi\left(l|\theta\right)$ and $v_{H}$ with probability $\pi\left(h|\theta\right)$. We have the following result without loss of generality by Bergemann and Morris (2016).
Lemma 1 (Direct Data Trade).
The trade data consists of two signals $\mathcal{S}=\left\{l,h\right\}$ with
$\mu_{h}$$\geq$ $\frac{v_{L}}{v_{H}}$ (6)
$\mu_{l}$$\leq$ $\frac{v_{L}}{v_{H}}.$ (7)
We call inequalities (6) and (7) the producer’s obedience constraints. Under the constraints, the product pricing is the producer’s best response to the direct data trade in which each signal is a price recommendation to the producer: pricing $v_{L}$ after signal $l$ and $v_{H}$ after signal $h$.
Plugging (2) into (6) (7), the obedience constraints for the producer to follow pricing recommendations are respectively
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-(1-\frac{v_{L}}{v_{H}})\mu_{0}\pi\left(l|H\right)$$\geq$ $\frac{v_{L}}{v_{H}}-\mu_{0}$ (8)
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-(1-\frac{v_{L}}{v_{H}})\mu_{0}\pi\left(l|H\right)$$\geq$ 00 0 (9)
Now we turn to the producer’s acceptance decision over the intermediary’s data trade offer ($\left(\left\{\mathcal{S},\pi\right\},T\right)$. When the producer does not accept the intermediary’s data offer, he has no more information than the prior $\mu_{0}$ and then the revenue is $R_{0}=\max\{v_{H}\mu_{0},v_{L}\}$. The producer accepts the data offer if and only if the product revenue from accepting is weakly larger than the one from rejecting
$R\left(\pi\right)-T\geq R_{0}$ (10)
where $R(\pi)$ is from (4).
In the following, we discuss the value of data to the producer under the direct data trade. Given bought data $\pi$, the producer charges $v_{L}$ (and all the consumers purchase the product) with a probability $\mu_{0}\pi(l|H)+(1-\mu_{0})\pi(l|L)$, charges $v_{H}$ (and only the type-H consumers will purchase the product) with a probability $\mu_{0}\pi(h|H)+(1-\mu_{0})\pi(h|L)$. Hence, we can rewrite the revenue (4) as
$R\left(\pi\right)$ $=$ $\left[\mu_{0}\pi\left(l|H\right)+\left(1-\mu_{0}\right)\pi\left(l|L\right)\right]v_{L}+\left[\mu_{0}\pi\left(h|H\right)+\left(1-\mu_{0}\right)\pi\left(h|L\right)\right]\mu_{h}v_{H}$
$=$ $\left[\mu_{0}\pi\left(l|H\right)+\left(1-\mu_{0}\right)\pi\left(l|L\right)\right]v_{L}+\mu_{0}\pi\left(h|H\right)v_{H}$
where the second equality follows (2).
The net gain from the data trade to the producer is then
$R\left(\pi\right)-R_{0}$ $=$ $\left[\mu_{0}\pi\left(l|H\right)+\left(1-\mu_{0}\right)\pi\left(l|L\right)\right]v_{L}+\mu_{0}\left(1-\pi\left(l|H\right)\right)v_{H}-\max\{v_{H}\mu_{0},v_{L}\}$ (11)
$=$ $-\mu_{0}\left(v_{H}-v_{L}\right)\underbrace{\left[\pi\left(l|H\right)-\mathbbm{1}_{\mu_{0}<\frac{v_{L}}{v_{H}}}\right]}_{\text{ surplus extraction }}+\left(1-\mu_{0}\right)v_{L}\underbrace{\left[\pi\left(l|L\right)-\mathbbm{1}_{\mu_{0}<\frac{v_{L}}{v_{H}}}\right]}_{\text{market inclusion}}$
where $\mathbbm{1}$ denotes the indicator function. [^13] The special case of $\mu_{0}=v_{L}/v_{H}$ is equivalent to either the case $\mu_{0}>v_{L}/v_{H}$ or $\mu_{0}<v_{L}/v_{H}$. The data trade has two effects on the producer’s revenue. The first is the change in the probability of serving the type-L consumers. With data trade the producer sells to the type-L consumers (of mass $\left(1-\mu_{0}\right)$) with probability $\pi\left(l|L\right)$. Instead, Without data trade the type-L consumers are served with probability $\mathbbm{1}_{\mu_{0}<\frac{v_{L}}{v_{H}}}$. This is the extensive market inclusion effect of data trade.
The second effect is the change in the probability of yielding surplus to the type-H consumers. With data trade the producer charges the type-H consumers (of mass $\mu_{0}$) a low price $v_{L}$ with probability $\pi\left(l|H\right)$. Without data trade the type-H consumers are charged the low price $v_{L}$ with probability $\mathbbm{1}_{\mu_{0}<\frac{v_{L}}{v_{H}}}$. This is the intensive surplus extraction effect of data trade. [^14] It is direct that when $\mu_{0}<\frac{v_{L}}{v_{H}}$, the market inclusion effect is negative: The market serving the type-L consumers shrinks. The surplus extraction effect is positive: The producer extracts the type-H consumer surplus with a larger probability. Instead, when $\mu_{0}>\frac{v_{L}}{v_{H}}$, the market inclusion effect is positive: The market serving the type-L consumers expands. the surplus extraction effect is negative: the producer extracts the type-H consumers surplus with a smaller probability.
The triangle gray set in Figure 3 illustrates the feasible data trade gain, which depends on the prior. Take $\mu_{0}>\frac{v_{L}}{v_{H}}$ as an example. Without data trade, the revenue is $\mu_{0}v_{H}$ which is at point $A$. It can go up to point $B$ under a data trade which induces posteriors $\mu_{l}$ and $\mu_{h}$. If the traded data is of full information, the intermediary extracts the maximum data trade surplus given $\mu_{0}$, which is $\mu_{0}v_{H}+\left(1-\mu_{0}\right)v_{L}$.
[…]
Figure 3: The triangle gray part is the feasible data trade gain from a data trade.
3.2 Consumers Optimality
Consumers of type $H$.
If the type-H consumers accept the service offer $\left(\mathcal{X},x,f\right)$ and report $\theta^{\prime}$ to the intermediary, the payoff in (5) is
$V-\left(x_{\theta^{\prime}}-H\right)^{2}-f_{\theta^{\prime}}+\left(v_{H}-v_{L}\right)\pi\left(l|\theta^{\prime}\right)$
where $\left(v_{H}-v_{L}\right)\pi\left(l|\theta^{\prime}\right)$ is the expected surplus for the type-H consumers in the product market. By reporting $\theta^{\prime}$, a type-H consumer will be charged $v_{L}$ and obtain surplus $v_{H}-v_{L}$ with probability $\pi\left(l|\theta^{\prime}\right)$. The data trade affects the probability of getting the surplus. The type-H consumers’ reporting affects both the service and its fee in the service market and the potential surplus in the product market. Hence, the type-$H$ consumers report the truth if and only if
$V-\left(x_{H}-H\right)^{2}-f_{H}+\left(v_{H}-v_{L}\right)\pi\left(l|H\right)\geq V-\left(x_{L}-H\right)^{2}-f_{L}+\left(v_{H}-v_{L}\right)\pi\left(l|L\right)$ (12)
If the type-H consumers reject the service offer, they will get nothing in the service market. In this case, the producer will treat consumers as coming from the population and make product pricing decisions based on the prior $\mu_{0}$. [^15] The intermediary cannot collect consumers data via providing service if the offer is rejected, and hence the producer cannot get extra information about consumers. That is, the product price $p_{0}$ without extra data is determined by (3) with the belief $\mu_{0}$. So the type-H consumers’ continuation value of rejecting the intermediary’s contract is
$\max\left\{v_{H}-p_{0},0\right\}\in\left\{\begin{array}[]{cc}\left\{0\right\}&\text{if }\mu_{0}>\frac{v_{L}}{v_{H}}\\ \{v_{H}-v_{L},0\}&\text{if }\mu_{0}=\frac{v_{L}}{v_{H}}\\ \left\{v_{H}-v_{L}\right\}&\text{if }\mu_{0}<\frac{v_{L}}{v_{H}}\end{array}\right.$
Hence, the type-H consumers accept the contract if and only if
$V-\left(x_{H}-H\right)^{2}-f_{H}+\left(v_{H}-v_{L}\right)\pi\left(l|H\right)\geq\max\left\{v_{H}-p_{0},0\right\}$ (13)
Consumers of type $L$.
The type-$L$ consumers get nothing in the service market if rejecting the intermediary’s service offer $\left(\mathcal{X},x,f\right)$. In the product market the producer will charge $p_{0}$ and the type-L consumers’ payoff is
$\max\left\{v_{L}-p_{0},0\right\}=0.$
If the type-L consumers accept the offer and reporting $\theta^{\prime}$, the type-L consumers receive $V-(x_{\theta^{\prime}}-L)^{2}-f_{\theta^{\prime}}$ in the service market and they always get zero in the product market as $\mathbb{E}_{\pi\left(s|\theta^{\prime}\right)}[\max\{v_{L}-p_{s},0\}]=0$. Thus, the type-L consumers’ payoff from accepting the offer and reporting $\theta^{\prime}$ in (5) is
$V-\left(x_{\theta^{\prime}}-L\right)^{2}-f_{\theta^{\prime}}$
Hence, the type-L consumers report the truth if and only if
$V-\left(x_{L}-L\right)^{2}-f_{L}\geq V-\left(x_{H}-L\right)^{2}-f_{H}$ (14)
and accept the intermediary’s contract if and only if
$V-\left(x_{L}-L\right)^{2}-f_{L}\geq 0$ (15)
3.3 Intermediary Optimality
The intermediary maximizes the total profit (1) from both the service and data markets by specifying a contract subject to the producer’s obedience conditions (8) (9) and participation constraint (10), the consumers’ incentive constraints (12) (14) and participation constraints (13) (15), and the following two feasibility conditions of data trade
$\pi\left(l|L\right)\text{,}\text{ }\pi\left(l|H\right)\in\left[0,1\right]$ (16)
As the intermediary is a monopolist in the data market, it extracts all the trade surplus by setting $T=R\left(\pi\right)-R_{0}$. The design setting is non-standard because of the presence of data trade, which endogenizes the consumer surplus.
We make the following short notations for simplicity.
Notation 1.
$\delta\equiv H-L$, $\Delta v\equiv v_{H}-v_{L}$, and $C\equiv(1-\mu_{0})v_{L}/\left(\mu_{0}\Delta v\right)$.
The parameter $\delta^{2}$ captures the variation in consumers’ tastes, reflecting the horizontal differentiation level in the service market, while $\Delta v$ captures the variation in consumers’ willingness to pay, measuring the vertical differentiation level in the product market.
It is ready to characterize the intermediary optimal mechanism. First, we assume a binary service space for better intuitions about the role of data trade in Section 4. Second, the section 5 extends the analysis to a general service case.
4 Binary Service
This section characterizes the intermediary optimal mechanism under the assumption of binary service $\mathcal{X}=\left\{L,H\right\}$ to focus on the cross-market role of data trade. [^16] One can treat the assumption as a delegated provision of service by the intermediary to its department of marketing, while the data trade is directly controlled by the intermediary. The marketing department’s choice of service level does not coordinate with the data trade policy. This is one of the practices: the intermediary is usually not the producer of service, but recommends service.
Lemma 2.
If $\mathcal{X}=\left\{L,H\right\}$, then $x_{H}=H$ and $x_{L}=L$.
Under the binary service, there is no distortion in the service provided to consumers in the optimal mechanism. When the service is binary, the service mismatch if any is fixed at the horizontal differentiation level $\delta^{2}$. Distorting a service provision requires a large compensation to consumers. Instead, without a mismatch in service provision, the intermediary can extract the most consumer surplus by charging a higher service fee.
The obedience constraint in (8) for charging a high product price, the producer’s revenue without any data in (10), and the type-H consumers’ participation constraint in (13) depend on the market prior $\mu_{0}$. The following discusses by two cases: $\mu_{0}>v_{L}/v_{H}$ and $\mu_{0}<v_{L}/v_{H}$. [^17] The special parameter case of $\mu_{0}=v_{L}/v_{H}$ is captured by either case.
4.1 Intermediary Optimality for $\mu_{0}>v_{L}/v_{H}$
When $\mu_{0}>\frac{v_{L}}{v_{H}}$, the obedience constraint for the producer to charge a high product price $OB_{H}$ (8) is redundant: whenever (9) holds, (8) is also true. To see this, recall $\frac{v_{L}}{v_{H}}-\mu_{0}<0$ and it makes (8) redundant given (9). In other words, if the producer follows the intermediary’s low price recommendation, then the producer would also follow the high price recommendation. We only need to consider $OB_{L}$ (9) for $\mu_{0}>\frac{v_{L}}{v_{H}}$. In particular, if (9) is binding, then $\pi\left(l|H\right)=C\cdot\pi\left(l|L\right)$. [^18] As $\mu_{0}>v_{L}/v_{H}$, $C\in\left(0,1\right)$ from Notation 1.
The type-H consumers will be still charged a product price $p_{0}=v_{H}$ even if they reject the service offer, for $\mu_{0}>\frac{v_{L}}{v_{H}}$. Hence, the type-H consumers’ outside option $\max\left\{v_{H}-p_{0},0\right\}$ is zero. The participation constraint of the type-H consumers (13) is
$V-\left(x_{H}-H\right)^{2}-f_{H}+\left(v_{H}-v_{L}\right)\pi\left(l|H\right)\geq 0$ (17)
When $\mu_{0}>\frac{v_{L}}{v_{H}}$, the producer charges $v_{H}$ without buying any data. The type-H consumers purchase the product and the type-L consumers are excluded from the market. The producer’s revenue is $R_{0}=\mu_{0}v_{H}$. The producer’s net gain from data trade in (11) becomes
$R\left(\pi\right)-R_{0}=\underbrace{-\mu_{0}\left(v_{H}-v_{L}\right)\pi\left(l|H\right)}_{\text{surplus extraction }}+\underbrace{\left(1-\mu_{0}\right)v_{L}\pi\left(l|L\right)}_{\text{market inclusion}}.$
The market inclusion effect of data trade is positive $(1-\mu_{0})\pi(l|L)>0$: The data trade extends the product market by allowing the producer to serve the type-L consumers (of mass $1-\mu_{0}$) with a higher probability $\pi(l|L)>0$. The surplus extraction effect of data trade is negative $-\mu_{0}\Delta v\pi(l|H)<0$. The producer extracts less type-H consumer surplus with data trade. With data trade, the producer charges the type-H consumers a low price $v_{L}$ with probability $\pi\left(l|H\right)$. Instead, without any data trade, all the type-H consumers are charged the high price $v_{H}$ and the intermediary does not yield any surplus to consumers.
The first main result on the intermediary’s optimal mechanism for data collecting and selling is summarized in the following proposition.
Proposition 1.
Suppose $\mu_{0}>v_{L}/v_{H}$. In the optimal mechanism, the intermediary sets the service fee to extract all the consumers surplus in both service and product markets
$f_{L}=V,\text{ }f_{H}=V+\pi\left(l|H\right)\Delta v$
The optimal data trade depends on the cross-market comparison between the mismatch cost in service $\delta^{2}$ and the vertical differentiation in the product $\Delta v$:
  1. 1.
    If $\delta^{2}\geq\Delta v$, the intermediary’s data trade is
    $\pi\left(l|L\right)=1,\text{ }\pi\left(l|H\right)\in\left[0,C\right]$
    and the data trade fee is
    $T=-\mu_{0}\pi\left(l|H\right)\Delta v+\left(1-\mu_{0}\right)v_{L}.$
    In this case, social welfare is
    $V+\left(1-\mu_{0}\right)v_{L}+\mu_{0}v_{H}.$
  2. 2.
    If $\delta^{2}<\Delta v$, the intermediary’s data trade is
    $\pi\left(l|L\right)=\frac{\delta^{2}}{\Delta v},\text{ }\pi\left(l|H\right)\in\Big{[}0,\frac{C\delta^{2}}{\Delta v}\Big{]}$
    and the data trade fee is
    $T=-\mu_{0}\pi\left(l|H\right)\Delta v+\left(1-\mu_{0}\right)v_{L}\frac{\delta^{2}}{\Delta v}.$
    In this case, social welfare is
    $V+\left(1-\mu_{0}\right)v_{L}\frac{\delta^{2}}{\Delta v}+\mu_{0}v_{H}.$
Irrelevance of $\pi\left(l|H\right)$ in intermediary’s objective.
The intermediary extracts all the surplus from both types of consumers in the service and product markets, the participation constraints of both types of consumers are binding. [^19] This is different from the standard mechanism design, where the type-H consumers have some surplus as information rent. With data trade, all consumers surplus are endogenous and extracted in the optimal mechanism. That is,
$f_{H}$ $=$ $V+\pi\left(l|H\right)\Delta v$
$f_{L}$ $=$ $V$
In the optimal mechanism, each type-H consumer has surplus $V$ from the service and gets surplus $\Delta v$ with probability $\pi\left(l|H\right)$ from the product. Each type-L consumer has a surplus $V$ from the service and nothing from the product.
However, in the data trade market, the intermediary needs to shift surplus $\Delta v$ with probability $\pi\left(l|H\right)$ to each type-H consumer. The intermediary in the data market works as the producer in the product market because it extracts all the data trade surplus. Due to the surplus-shifting, the probability of charging type-H consumers a low price $\pi\left(l|H\right)$ will not affect the intermediary’s payoff. To see this, plugging the binding $IR_{L}$ (15) and $IR_{H}$ (17): $f_{L}=V$, $f_{H}=V+\pi\left(l|H\right)\Delta v$ into the intermediary’s objective function
$V+\mu_{0}\underbrace{\pi\left(l|H\right)\Delta v}_{\text{surplus-extracting}}-\mu_{0}\underbrace{\pi\left(l|H\right)\Delta v}_{\text{surplus-shifting }}+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}$
where the profit from the service market is $V+\mu_{0}\pi\left(l|H\right)\Delta v$ and from the data market is $-\mu_{0}\pi\left(l|H\right)\Delta v+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}$. The surplus-shifting and the surplus-extracting offset each other. [^20] This is true because the intermediary has full bargaining power over the data trade surplus as it is a monopolist in the data market.
Intermediary’s tradeoff.
With binding $IR_{L}$ (15) and $IR_{H}$ (17), then $IC_{L}$ (14) will be redundant, the intermediary only needs to consider $OB_{L}$ (9) and $IC_{H}$ (12). The intermediary’s reduced problem is to choose $\pi$ to solve
$\max_{\pi}$ $V+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}$
subject to $\pi\left(l|L\right)\Delta v-\delta^{2}\leq 0$
$\pi\left(l|H\right)\leq C\pi\left(l|L\right)$
In the first constraint: $\pi\left(l|L\right)\Delta v-\delta^{2}\leq 0$, if a type-H consumer reports the truth, she will get nothing. Instead, if she misreports, she will get the surplus $\Delta v$ with probability $\pi\left(l|L\right)$ but incur mismatch loss in the service $\delta^{2}$. To incentivize the type-Hs to tell the truth, the intermediary’s choice of $\pi\left(l|L\right)$ has to satisfy $\pi\left(l|L\right)\Delta v\leq\delta^{2}$.
This reduced problem clarifies the tradeoff faced by the intermediary. The type-H consumers’ expected surplus in the product by misreporting, $\pi\left(l|L\right)\Delta v$ is bounded by the mismatch in the service $\delta^{2}$. However, the intermediary prefers $\pi\left(l|L\right)$ as high as possible because it can extract the type-L consumers’ surplus with a higher probability in the product market.
For the intermediary, there are two instruments to incentivize the type-H consumers to tell the truth: decrease the probability of charging $v_{L}$ for those who report $L$, and the mismatch (wrong service) itself as a penalty. The mismatch is preferred by the intermediary because it does not require to lower $\pi\left(l|L\right)$. Hence, when the mismatch cost in service is high enough as a penalty, the type-H consumers will not misreport and the intermediary would like to set $\pi\left(l|L\right)$ as high as possible, that is $\pi\left(l|L\right)=1$. The type-L consumers are all served in the product market. In this case, the total profit of the intermediary is $V+\left(1-\mu_{0}\right)v_{L}$. Social welfare is the sum of consumers surplus 00, producer surplus $\mu_{0}v_{H}$ and intermediary surplus $V+\left(1-\mu_{0}\right)v_{L}$, which is $V+\left(1-\mu_{0}\right)v_{L}+\mu_{0}v_{H}$.
When the mismatch cost in service is low as a penalty for the type-H consumers’ misreport $\delta^{2}<\Delta v$, the mismatch itself cannot provide sufficient incentive for the type-H consumers to reveal their information. Instead, the intermediary would like to lower $\pi\left(l|L\right)$ to incentivize truth-telling to collect information. In this case, some type-L consumers are excluded from the product market. The intermediary’s optimal profit from the service market is $V+\mu_{0}\pi\left(l|H\right)\Delta v$ and profit from the data trade is $-\mu_{0}\pi\left(l|H\right)\Delta v+\left(1-\mu_{0}\right)v_{L}\frac{\delta^{2}}{\Delta v}$. The equilibrium total payoff to the intermediary is $V+\left(1-\mu_{0}\right)v_{L}\frac{\delta^{2}}{\Delta v}$ and social surplus is $V+\left(1-\mu_{0}\right)v_{L}\frac{\delta^{2}}{\Delta v}+\mu_{0}v_{H}$.
Following Proposition 1, we have the following results.
Corollary 1.
Suppose $\mu_{0}>v_{L}/v_{H}$. No information data trade cannot be optimal.
To see it by contradiction. Suppose in the optimal mechanism $\pi\left(l|H\right)=\pi\left(l|L\right)$. In either case of $\delta^{2}\geq\Delta v$ or $\delta^{2}<\Delta v$, the obedience constraint (9) requires $\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}<\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)$ which contradicts with $\mu_{0}>\frac{v_{L}}{v_{H}}$. Intuitively, when $\mu_{0}>\frac{v_{L}}{v_{H}}$ any extra information will make it easier for the producer to price discriminate the consumers. This is because in the case of $\mu_{0}>\frac{v_{L}}{v_{H}}$ the market inclusion effect is positive: without any data trade all the type-L consumers are excluded in the product market.
Corollary 1 works as a rationale for the so-called digital “privacy paradox”: the general discrepancies between consumers’ active data-sharing behaviors and their self-stated privacy concerns. When $\mu_{0}>\frac{v_{L}}{v_{H}}$, trading some (even full) data is optimal. In other words, the highest privacy protection is not preferred. The results rationalize the paradox as a market outcome. Consumers would prefer to reveal some (even full) information about their private preferences even though they are informed that their revealed information might be traded and used to price against them in other markets. This is possible when the benefits of better service outweigh the cost of losing surpluses in the product. So the intermediary does not need to reduce probability $\pi\left(l|L\right)$. The choice of $\pi\left(l|H\right)$ also makes the producer obedient.
4.2 Intermediary Optimality for $\mu_{0}<v_{L}/v_{H}$
When $\mu_{0}<\frac{v_{L}}{v_{H}}$, $\max\{v_{H}-p_{0},0\}=\Delta v$ and the obedience constraint for the producer to charge a low price in (9) is redundant. The participation constraint of the type-H consumers (13) is [^21] In the special parameter case of $\mu_{0}=v_{L}/v_{H}$, $p_{0}$ equals to $v_{H}$ or $v_{L}$ and then (13) is either (17) or (18).
$V-\left(x_{H}-H\right)^{2}-f_{H}+\left(v_{H}-v_{L}\right)\pi\left(l|H\right)\geq\Delta v$ (18)
Without data trade the producer will charge $v_{L}$ and all consumers purchase in the product market. The producer’s revenue is $R_{0}=v_{L}$ without data trade. The net gain in the product revenue from data trade (11) becomes
$R\left(\pi\right)-R_{0}=-\mu_{0}\left[\pi\left(l|H\right)-1\right]\left(v_{H}-v_{L}\right)+\left(1-\mu_{0}\right)\left[\pi\left(l|L\right)-1\right]v_{L}$
Here, the market inclusion effect of data trade is negative: With data trade, the market serving the type-L consumers shrinks from $1$ to $\pi\left(l|L\right)$. The surplus extraction effect of data trade is positive: The intermediary extracts extra surplus from the type-H consumers by $\left[1-\pi\left(l|H\right)\right]\left(v_{H}-v_{L}\right)$. As we will see, the extra extracted surplus due to data trade will be shifted back to the consumers in the service market.
Proposition 2.
Suppose $\mu_{0}<v_{L}/v_{H}$. In the optimal mechanism, the intermediary sets service fees to extract all the consumers surplus in the service and product markets
$f_{L}=V,\text{ }f_{H}=V-\left(1-\pi\left(l|H\right)\right)\Delta v$
and the optimal data trade depends on the cross-market comparison between the mismatch cost $\delta^{2}$ and the vertical differentiation $\Delta v$
$\pi\left(l|L\right)=1,\text{ }\pi\left(l|H\right)\in\left[\max\left\{1-\frac{\delta^{2}}{\Delta v},0\right\},1\right].$
The data trade fee
$T=\mu_{0}\Delta v\left(1-\pi\left(l|H\right)\right).$
and social welfare is
$V+v_{L}+\mu_{0}\Delta v.$
Similarly, $\pi\left(l|H\right)$ is canceled out in the intermediary’s objective due to the offset of surplus-shifting in the product and surplus-extracting in the service. By plugging the binding $IR_{L}$ (15) and $IR_{H}$ (18): $f_{L}=V$, $f_{H}=V+\left[\pi\left(l|H\right)-1\right]\Delta v$ into the objective, the intermediary solves the reduced problem subject to $IC_{L}$ (14) and $OB_{H}$ (8)
$\max_{\pi}$ $V+\left(1-\mu_{0}\right)v_{L}\pi\left(l|L\right)-\left(1-\mu_{0}\right)v_{L}$
subject to $\left(1-\pi\left(l|H\right)\right)\Delta v-\delta^{2}\leq 0$
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)\geq\frac{v_{L}}{v_{H}}-\mu_{0}$
The intermediary would set $\pi\left(l|L\right)=1$ as there is no constraint on it. At $\pi\left(l|L\right)=1$, any $\pi\left(l|H\right)\in\left[0,1\right]$ satisfies $OB_{H}$ (8). But $IC_{L}$ (14) imposes the restrictions on the choice set of $\pi\left(l|H\right)$, which relies on the mismatch as a penalty for misreporting. When the mismatch is high $\delta^{2}>\Delta v$, the penalty of service mismatch itself can incentivize type-L consumers to report the truth. So the intermediary can choose any $1-\pi\left(l|H\right)\in\left[0,1\right]$. In this case, the service profit is $V-\mu_{0}\Delta v\left[1-\pi\left(l|H\right)\right]$ and the data trade profit is $\mu_{0}\Delta v\left[1-\pi\left(l|H\right)\right]$. So its optimal total payoff is $V$.
When the mismatch is low $\delta^{2}<\Delta v$, the penalty of service mismatch itself cannot provide strong incentives for type-L consumers to tell the truth. The intermediary chooses $1-\pi\left(l|H\right)\in\big{[}0,\frac{\delta^{2}}{\Delta v}\big{]}$ such that the truth is induced. In this case, the optimal service profit is $V-\mu_{0}\left(1-\pi\left(l|H\right)\right)\Delta v$ and the data trade profit is $\mu_{0}\left(1-\pi\left(l|H\right)\right)\Delta v$. The intermediary’s total payoff is $V$.
In both cases, social welfare is the sum of intermediary surplus $V$, producer surplus $v_{L}$, and type-H consumers surplus $\mu_{0}\Delta v$:
$V+v_{L}+\mu_{0}\Delta v.$
As $\pi\left(l|L\right)=\pi\left(l|H\right)=1$ is optimal when $\mu_{0}<v_{L}/v_{H}$ from Proposition 2, it follows
Corollary 2.
Suppose $\mu_{0}<v_{L}/v_{H}$. The no-information data trade is always optimal.
The optimal data trade can disclose no information about consumers. Since the product market has included the type-L consumers when $\mu_{0}<\frac{v_{L}}{v_{H}}$, there is no space for the producer to extract type-L consumer surplus from the extra information of data trade. Hence, there is no price discrimination in the product market.
4.3 Optimal Data Trade
To be clear on the tradeoff of data trade, we summarize the results on data trade in Table 1. Several implications follow the optimal data trade results.
Table 1: Optimal Data Trade
Prior Mismatch Data Trade
$\mu_{0}>\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$ $\pi\left(l|L\right)=1,\pi\left(l|H\right)\in\left[0,C\right]$
$\delta^{2}<\Delta v$ $\pi\left(l|L\right)=\frac{\delta^{2}}{\Delta v},\pi\left(l|H\right)\in\left[0,C\frac{\delta^{2}}{\Delta v}\right]$
$\mu_{0}<\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$ $\pi\left(l|L\right)=1,\pi\left(l|H\right)\in\left[0,1\right]$
$\delta^{2}<\Delta v$ $\pi\left(l|L\right)=1,\pi\left(l|H\right)\in\left[1-\frac{\delta^{2}}{\Delta v},1\right]$
First, the intermediary will not sell full information in the optimal mechanism when $\delta^{2}<\Delta v$222222 The result is mainly from the cross-market effect of data, different from the ratchet effect. To see this, take the case of $\mu_{0}>\frac{v_{L}}{v_{H}}$ as an example. In the optimal mechanism, the service profit is $V+\mu_{0}\pi\left(l|H\right)\Delta v$ and the data trade fee is $-\mu_{0}\pi\left(l|H\right)\Delta v+\left(1-\mu_{0}\right)v_{L}\frac{\delta^{2}}{\Delta v}$. In contrast, if the data trade is full information, the service profit decreases to $V+\mu_{0}\left(\delta^{2}-\Delta v\right)$, while the data trade fee increases to $\left(1-\mu_{0}\right)v_{L}\frac{\delta^{2}}{\Delta v}$. The loss in service profit of the amount $\mu_{0}\pi\left(l|H\right)\Delta v-\mu_{0}\left(\delta^{2}-\Delta v\right)$ is larger than the amount of increment in the data profit $\mu_{0}\pi\left(l|H\right)\Delta v$ when $\delta^{2}<\Delta v$. To preserve the service profit, the intermediary would prefer to sell only partial or no information. [^23] The total payoff from the full information $V+\mu_{0}\left(\delta^{2}-\Delta v\right)+\left(1-\mu_{0}\right)v_{L}$ is smaller than the one from the optimal data trade $V+\left(1-\mu_{0}\right)v_{L}\frac{\delta^{2}}{\Delta}$ when $\delta^{2}<\Delta v$. Instead, when $\delta^{2}\geq\Delta v$ selling full information would be optimal for the intermediary. The following result directly follows.
Corollary 3.
When $\delta^{2}<\Delta v$, the full information data trade cannot be optimal.
Second, the privacy protection laws should be careful about the suit levels of protection depending on the market environment. The privacy protection laws usually work as the intermediary’s commitment on how much data to trade. The highest privacy protection is sometimes optimal (e.g. small $\mu_{0}$), while trading full data is sometimes optimal (e.g., large $\delta^{2}$).
4.4 Welfare and Privacy Protection
The baseline framework with a holistic viewpoint provides an opportunity to evaluate the effects of the laws of privacy protection comprehensively: Taking into account both data collecting and selling. The booming digital economy featuring data trade and transfer brings wide concerns about privacy protections, which attracts attention from both the public and regulatory authorities. Several laws have been established across different countries. Despite the importance of privacy protections, a comprehensive understanding of the potential effects of regulations on the data market is still missing.
When the laws ban the data trade to pursue the highest level of consumer privacy protection, the service market, and product market work separately because they are not linked by the intermediary’s data trade. The data shared with the intermediary will no longer be used in another market against the consumers’ interests. In this environment, the type-L consumers’ total surplus is zero. The type-H consumers get some surplus when the product price is $v_{L}$, which corresponds to the case of $\mu_{0}<\frac{v_{L}}{v_{H}}$. In the product market, the producer extracts all the surplus under the prior $\mu_{0}$. The producer’s surplus is $R_{0}=\max\left\{v_{L},v_{H}\mu_{0}\right\}$.
Proposition 3.
Suppose there exists no data trade.
  1. 1.
    In the service market, the intermediary sets $x_{L}=L$ and $x_{H}=H$ and the service fees are $f_{L}=f_{H}=V$. Both types of consumers surplus in service are zero and the intermediary surplus is $V$.
  2. 2.
    In the product market, the producer sets price $p_{0}$. The type-L consumers surplus is 00. The type-H consumers surplus is $\Delta v$ when $\mu_{0}<\frac{v_{L}}{v_{H}}$ and is 00 when $\mu_{0}>\frac{v_{L}}{v_{H}}$. The producer gets $\max\left\{v_{L},\mu_{0}v_{H}\right\}$.
Without data trade, social welfare is $V+\mu_{0}v_{H}$ when $\mu_{0}>\frac{v_{L}}{v_{H}}$, and $V+v_{L}+\mu_{0}\Delta v$ when $\mu_{0}<\frac{v_{L}}{v_{H}}$.
To investigate the effects of banning data trade, we compare the surplus distribution and welfare with and without the data market.
Table 2: Surplus Distribution
With Data Market Without Data Market Prior Mismatch Intermediary Producer Consumer $H$ Consumer $L$ Intermediary Producer Consumer $H$ Consumer $L$ $\mu_{0}>\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$ $V+\left(1-\mu_{0}\right)v_{L}$ $\mu_{0}v_{H}$ 0 0 $V$ $\mu_{0}v_{H}$ 0 0 $\delta^{2}<\Delta v$ $V+\left(1-\mu_{0}\right)v_{L}\Delta v$ $\mu_{0}v_{H}$ 0 0 $V$ $\mu_{0}v_{H}$ 0 0 $\mu_{0}<\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$ $V$ $v_{L}$ $\Delta v$ 00 $V$ $v_{L}$ $\Delta v$ 00 $\delta^{2}<\Delta v$ $V$ $v_{L}$ $\Delta v$ 00 $V$ $v_{L}$ $\Delta v$ 00
Table 3: Surplus Change Led by Data Trade Ban
Prior Mismatch Intermediary Producer Consumer $H$ Consumer $L$
$\mu_{0}>\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$downarrow -- -- --
$\delta^{2}<\Delta v$downarrow -- -- --
$\mu_{0}<\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$- -- -- --
$\delta^{2}<\Delta v$- -- -- --
Surplus.
In Table 2, we summarize the consumers, the producer, and the intermediary surplus with and without data trade ban. The data trade ban does not affect the producer surplus and consumers surplus. The producer surplus is always the producer’s revenue in the product market without buying data, $R_{0}$. This is because the intermediary has the full bargaining power over the data trade surplus and the producer does not get any share of the data trade surplus even when there is a data market. The type-L consumers surplus is always zero: The intermediary extracts their surplus in the service market and they don’t have any surplus in the product market. The type-H consumers surplus is $\Delta v$ in the product market.
Banning data trade affects the intermediary only when $\mu_{0}>\frac{v_{L}}{v_{H}}$. The intermediary can only extract the service surplus $V$ without the data market. However, the intermediary gets better off by serving an extra proportion of type-L consumers due to the positive market inclusion effect of data trade. With data trade, the producer is able to price discriminate and extract more consumers surplus by $\left(1-\mu_{0}\right)v_{L}\min\left\{\Delta v,1\right\}$. Changes in surplus are summarized in Table 3.
Social Welfare.
Table 4 summarizes social welfare with and without data trade market and Table 5 shows changes in welfare. Banning data market decreases social welfare when $\mu_{0}>\frac{v_{L}}{v_{H}}$. Intuitively, without data trade, it becomes harder for the producer to price discriminate consumers and extract the most surplus. Specifically, the data trade brings two effects: surplus extraction and market inclusion. The surplus extraction effect does not matter for social welfare as it is a reallocation of surplus between the producer and consumers. The market inclusion effect of data trade is $\left(1-\mu_{0}\right)v_{L}\min\left\{\Delta v,1\right\}$ when the proportion of the type-L consumers is large. This positive effect will disappear if laws ban the data market.
Proposition 4.
Banning data trade decreases social welfare if the proportion of the type-H consumers in the market is large $\mu_{0}>v_{L}/v_{H}$.
Table 4: Welfare Summary
Prior Mismatch With Data Market Without Data Market
$\mu_{0}>\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$ $V+\left(1-\mu_{0}\right)v_{L}+\mu_{0}v_{H}$ $V+\mu_{0}v_{H}$
$\delta^{2}<\Delta v$ $V+\left(1-\mu_{0}\right)v_{L}\Delta v+\mu_{0}v_{H}$ $V+\mu_{0}v_{H}$
$\mu_{0}<\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$ $V+v_{L}+\mu_{0}\Delta v$ $V+v_{L}+\mu_{0}\Delta v$
$\delta^{2}<\Delta v$ $V+v_{L}+\mu_{0}\Delta v$ $V+v_{L}+\mu_{0}\Delta v$
Table 5: Change in Welfare Led by Ban
Prior Mismatch Welfare Change
$\mu_{0}>\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$downarrow
$\delta^{2}<\Delta v$downarrow
$\mu_{0}<\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$-
$\delta^{2}<\Delta v$-
5 General Service
The previous baseline analysis is under the binary service. As the mismatch level in service changes consumers’ incentives, this section investigates a general service space $\mathcal{X}=\mathbb{R}$ which allows flexible mismatch level. The analysis shows that the main insights hold with a general service. The general service generates a different cutoff for selling full information to be optimal.
The results rely on the market prior similar to the baseline. When $\mu_{0}<v_{L}/v_{H}$, there is no distortion in the intermediary optimal service provision even under a general service space. The results on service fee, data trade, and data fee are then as Proposition 2: selling no information can be optimal. We relegate them and their proofs to Appendix A.8 of Proposition 7.
We summarize the results for $\mu_{0}>v_{L}/v_{H}$ as follows.
Proposition 5.
Suppose the service space $\mathcal{X}=\mathbb{R}$ and $\mu_{0}>v_{L}/v_{H}$. In the optimal mechanisms, all the consumers surpluses are extracted. For $\delta^{2}\geq\Delta v$, there is no distortion in service provisions $x_{L}=L$, $x_{H}=H$, and the service fees, data trade are as the part 1 of Proposition 1.
For $\delta^{2}<\Delta v$, there is a downward distortion in service provided to type-$L$ consumers, but there is no service distortion for type-$H$ consumers $x_{H}=H$. The following hold:
  1. 1.
    If $\frac{\Delta v}{v_{L}+v_{H}}<\frac{\delta^{2}}{\Delta v}<1$, then the service for type-$L$ is $x_{L}=L-\frac{\Delta v-\delta^{2}}{2\delta}$, the service fees are $f_{L}=V-(\frac{\Delta v-\delta^{2}}{2\delta})^{2}$, $f_{H}=V+\pi(l|H)\Delta v$; and the data trade is
    $\pi(l|L)=1,\text{ }\pi(l|H)\in[0,C]$
    the data trade fee is
    $T=-\mu_{0}\pi(l|H)\Delta v+(1-\mu_{0})v_{L}$
    and the intermediary’s total revenue is
    $V-(1-\mu_{0})\left(\frac{\Delta v-\delta^{2}}{2\delta}\right)^{2}+(1-\mu_{0})v_{L}$
    In this case social welfare is
    $V-(1-\mu_{0})\left(\frac{\Delta v-\delta^{2}}{2\delta}\right)^{2}+(1-\mu_{0})v_{L}+\mu_{0}v_{H}$
  2. 2.
    If $\frac{\delta^{2}}{\Delta v}<\frac{\Delta v}{v_{L}+v_{H}}$, then the service for type-$L$ is $x_{L}=L-\frac{\delta}{\Delta v}v_{L}$, the service fees are $f_{L}=V-(\frac{\delta}{\Delta v}v_{L})^{2}$, $f_{H}=V+\pi(l|H)\Delta v$; and the data trade is
    $\pi(l|L)=\frac{\delta^{2}}{\Delta v}\frac{v_{H}+v_{L}}{\Delta v},\text{ }\pi(l|H)\in\Big{[}0,C\frac{\delta^{2}}{\Delta v}\frac{v_{H}+v_{L}}{\Delta v}\Big{]}$
    the data trade fee is
    $T=-\mu_{0}\pi(l|H)\Delta v+(1-\mu_{0})v_{L}\frac{\delta^{2}}{(\Delta v)^{2}}(v_{L}+v_{H})$
    and the intermediary’s total revenue is
    $V+(1-\mu_{0})\left(\frac{\delta}{\Delta v}\right)^{2}v_{L}v_{H}.$
    In this case social welfare is
    $V+(1-\mu_{0})\left(\frac{\delta}{\Delta v}\right)^{2}v_{L}v_{H}+\mu_{0}v_{H}.$
When $\mu_{0}>v_{L}/v_{H}$ and the mismatch cost is large enough, i.e., $\delta^{2}>\Delta v$, there is no distortion in consumers’ service even with a general service space. This is because the mismatch cost in service is so high that consumers are willing to report the private information without any further distortion in service level. The intermediary sets the highest $\pi\left(l|L\right)=1$ and does not need to distort any service level to maintain the type-H’s incentive. The results on service fees and data trade are as the part 1 of Proposition 1.
However, when $\mu_{0}>v_{L}/v_{H}$ and the mismatch cost is small, $\delta^{2}<\Delta v$, the intermediary needs to downward distort the type-L’s service. In this case, the penalty of the wrong service cannot provide enough incentive. The intermediary will manipulate the type-H consumers’ incentive via other two instruments: the probability of surplus in the product and service provision. Although distorting the service will reduce the surplus extracted from the type-Ls in the service market, the intermediary can extract more surplus from the type-H consumers in the product market. When the benefit of more surplus from type-H consumers outweighs the cost of reduced surplus from type-L consumers, the intermediary would prefer to distort the service to the type-L consumers. This happens when the population of type-H consumers is large enough $\mu_{0}>v_{L}/v_{H}$.
In sum, the results on data trade are similar to the baseline analysis, but with a different threshold for selling full information to be optimal. Under a general service space selling full information can be optimal when the mismatch cost is sufficiently large, i.e., $\delta^{2}>(\Delta v)^{2}/(v_{L}+v_{H})$ by the part 2 of Proposition 5. The cutoff $(\Delta v)^{2}/(v_{L}+v_{H})$ is smaller than the one, $\Delta v$ in baseline results under binary service.
As both service distortion and price discrimination affect consumers’ incentives, the general service setup allows us to discuss their relations in the optimal mechanisms. The below follows directly from propositions 5 and 7.
Corollary 4 (Service distortion and price discrimination).
In the intermediary optimal mechanisms, the service distortion and no-price discrimination are exclusive.
Given a service distortion, we have $\mu_{0}>v_{L}/v_{H}$ and $\delta^{2}<\Delta v$; otherwise, there is no distortion by Proposition 5 and 7. From $\mu_{0}>v_{L}/v_{H}$ and $\delta^{2}<\Delta v$, it is impossible that $\pi(l|L)=\pi(l|H)$, which implies no-discrimination is not possible.
Instead, given no price discrimination, it has to be $\mu_{0}<v_{L}/v_{H}$ by contradiction: Suppose $\mu_{0}>v_{L}/v_{H}$, then it is impossible for $\pi(l|H)=\pi(l|L)$. Hence, $\mu_{0}<v_{L}/v_{H}$ implies that there is no service distortion from Proposition 7.
Regulation.
Similarly, we discuss the welfare implications of banning data trade within this more general setup. The main welfare implications are same with those in the binary service and are summarized into the following proposition.
Proposition 6.
Suppose a general service space. Banning data trade will decrease social welfare for $\mu_{0}>v_{L}/v_{H}$, and will not affect social welfare for $\mu_{0}<v_{L}/v_{H}$.
6 Discussion
Competition between Data Buyers.
In the stylized model, there is one data buyer, i.e., the producer. When there are two producers and both of them buy data from the intermediary, the pricing competition in the product market between two producers might compete away the producer’s profit and hence prevent the intermediary from a beneficial data trade. In this sense, the intermediary would prefer an exclusive data trade, selling data only to one producer, which goes back to the baseline framework of one data buyer.
Commitment of the Intermediary.
Recall that our work focuses on the cross-market incentive instead of the inter-temporal one. If one interprets our results from the inter-temporal viewpoint, there can be a commitment problem of the intermediary under some conditions. Take the result in Corollary 3 as an example, when $\delta^{2}<\Delta v$, the data trade ex ante committed by the optimal mechanism can be inconsistent with what will happen after consumers report the true types in the service market. The optimal mechanism will never sell full information when $\delta^{2}<\Delta v$ while the intermediary has incentives to sell all the information to get the maximum data trade gain. [^24] The intermediary’s commitment is not a problem when $\delta^{2}$ is large enough. The full information data trade is optimal if only if $\delta^{2}$ is sufficiently large. Given a full information data trade in the optimality, there will not be the time-inconsistent behaviors of the intermediary after collecting the data. Bester and Strausz (2001), Doval and Skreta (2022) and Doval and Skreta (Forthcoming) study the mechanism with the principal’s limited commitment.
7 Conclusion
This paper develops an integrated framework of the intermediary’s data sourcing from consumers and data trade to a third party: Consumers benefit from revealing information about their taste to the intermediary for better-matched service. A third party buys data from the intermediary so that it can extract more surplus from consumers in the product market. We characterize the optimal mechanisms for collecting and selling data. Due to the cross-market role of data, the equilibrium data trade depends on the consumers’ horizontal differentiation level in the data-sourcing service market and the vertical one in the product market.
In the intermediary optimal mechanism, the intermediary sells no information or partial information to preserve profits in the service market when consumers’ mismatch cost from a wrong service is low. The low mismatch cost itself cannot sufficiently incentivize consumers to reveal information. The intermediary has to compensate consumers more (by decreasing service fees) to sell full information. However, the loss in service profit outweighs the net gain from selling full information rather than partial or no information. Hence, the intermediary would prefer to sell partial or no information, which suggests that the market can protect consumer privacy without regulations. In contrast, the intermediary can sell full information to data buyers when consumers’ mismatch cost in service is high. In this case, consumers’ privacy protection is at the lowest level.
Using the framework to evaluate privacy protection, we find that a ban on data trade may reduce social welfare because it makes it harder to price discriminate in the product market. The finding suggests an alternative policy to regulate the data market, such as a lump-sum tax. In this way, there will not be welfare loss, and consumers and the producer can share some surplus with the intermediary via government transfers.
Appendix A Appendix
A.1 Proof of Lemma 1
Proof.
The basic idea is to merge the signals that induce the same posterior (Bergemann and Morris, 2016). Let $\hat{\pi}(l|\theta)=\sum\limits_{s:\mu_{s}<v_{L}/v_{H}}\pi(s|\theta)$ and $\hat{\pi}(h|\theta)=\sum\limits_{s:\mu_{s}\geq v_{L}/v_{H}}\pi(s|\theta)$. Then, the data trade is captured by ($(\{l,h\},\hat{\pi})$. In the content after Lemma 1, we use the notation of ($(\{l,h\},\pi)$ for the binary direct data trade ($(\{l,h\},\hat{\pi})$ without confusions. ∎
A.2 Proof of Lemma 2
For a general service space $\mathcal{X}$, we have results in the following Lemma 3. Lemma 2 directly follows Lemma 3 with $\mathcal{X}=\{L,H\}$.
Lemma 3.
Suppose a general service space $\mathcal{X}$. In the intermediary optimal mechanism, the service provided to type-H consumers $x_{H}\geq H$ and to type-L consumers $x_{L}\leq L$.
Proof.
To make the proof easy to read, we write out repeatedly the intermediary’s problem in the general space $\mathcal{X}$. That is, $\max\limits_{x,f,\pi,T}f_{H}\mu_{0}+f_{L}\left(1-\mu_{0}\right)+T$ subject to
$V-\left(x_{H}-H\right)^{2}-f_{H}+\pi\left(l|H\right)\Delta v\geq V-\left(x_{L}-H\right)^{2}-f_{L}+\pi\left(l|L\right)\Delta v$ ($IC_{H}$ )
$V-\left(x_{L}-L\right)^{2}-f_{L}\geq V-\left(x_{H}-L\right)^{2}-f_{H}$ ($IC_{L}$)
$V-\left(x_{H}-H\right)^{2}-f_{H}+\pi\left(l|H\right)\Delta v\geq\max\left\{v_{H}-p_{0},0\right\}$ ($IR_{H}$)
$V-\left(x_{L}-L\right)^{2}-f_{L}\geq 0$ ($IR_{L}$)
$\mu_{0}\left(1-\pi\left(l|H\right)\right)\geq\frac{v_{L}}{v_{H}}\left[\mu_{0}\left(1-\pi\left(l|H\right)\right)+\left(1-\mu_{0}\right)\left(1-\pi\left(l|L\right)\right)\right]$ ($OB_{H}$)
$\frac{v_{L}}{v_{H}}\left[\mu_{0}\pi\left(l|H\right)+\left(1-\mu_{0}\right)\pi\left(l|L\right)\right]\geq\mu_{0}\pi\left(l|H\right)$ ($OB_{L}$)
$R\left(\pi\right)-R_{0}\geq T$ ($IR_{S}$)
$\pi\left(l|L\right),\pi\left(l|H\right)\in\left[0,1\right]$ ($FE$)
First, $x_{H}\geq H$ in the optimal mechanism. To see this by contradiction. Suppose $x_{H}<H$, then $\left(x_{H}-H\right)^{2}>0$. Increasing $f_{H}$ while decreasing $\left(x_{H}-H\right)^{2}$ by equal size will not affect $IC_{H}$ (12) $IR_{H}$ (13) and $IR_{L}$ (15). This change does not affect $IC_{L}$ (14) either. To see this, there are two cases to consider: $x_{H}<L$ and $L<x_{H}<H$. Case 1: when $x_{H}<L$, decreasing $\left(x_{H}-H\right)^{2}$ means increasing $x_{H}$, which will decrease
$V-\left(x_{H}-L\right)^{2}-f_{H}=V-\left(x_{H}-H\right)^{2}-f_{H}-\delta^{2}-2\delta\left(x_{H}-H\right),$
because changes in $-\left(x_{H}-H\right)^{2}$ and $-f_{H}$ are canceled out, and $-2\delta\left(x_{H}-H\right)$ decreases. So $IC_{L}$ (14) is not affected. Case 2: when $L<x_{H}<H$, decreasing $\left(x_{H}-H\right)^{2}$ means increasing $x_{H}$, which will increase $\left(x_{H}-L\right)^{2}$. Then $IC_{L}$ (14) obviously holds in this case.
Thus, increasing $f_{H}$ while decreasing $\left(x_{H}-H\right)^{2}$ by equal size will not affect $IC_{L}$ (14) and other constraints while increasing the intermediary’s payoff. Hence, $x_{H}<H$ is not optimal. In the optimality, it must be $x_{H}\geq H$.
Second, in the optimal mechanism $x_{L}\leq L$. Suppose $x_{L}>L$, then $\left(x_{L}-L\right)^{2}>0$. Increasing $f_{L}$ while decreasing $\left(x_{L}-L\right)^{2}$ by equal size will not affect $IC_{L}$ (14) $IR_{H}$ (13) and $IR_{L}$ (15). It does not affect $IC_{H}$ (12) either. To see this, decreasing $\left(x_{L}-L\right)^{2}$ means decreasing $x_{L}$ which will decrease
$V-\left(x_{L}-H\right)^{2}-f_{L}=V-\left(x_{L}-L\right)^{2}-f_{L}-\delta^{2}-2\delta\left(x_{L}-L\right)$
because changes in $-\left(x_{L}-L\right)^{2}$ and $-f_{L}$ are canceled out, and $-2\delta\left(x_{L}-L\right)$ decreases as $x_{L}$ decreases. So increasing $f_{L}$ while decreasing $\left(x_{L}-L\right)^{2}$ by equal size will not affect $IC_{H}$ (12) and thus not affect all the constraints but increase the intermediary’s payoff. Hence, $x_{L}>L$ cannot be optimal.
In particular, with $\mathcal{X}=\{H,L\}$, $x_{L}=L$ and $x_{H}=H$ which proves Lemma 2.
A.3 Proof of Proposition 1
Proof.
When $\mu_{0}>\frac{v_{L}}{v_{H}}$. From $OB_{H}$ (8) and $OB_{L}$ (9)
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)$$\geq$ $\frac{v_{L}}{v_{H}}-\mu_{0}$
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)$$\geq$ 00 0
The obedience constraint for the producer to charge a high price (8) is redundant, we can ignore (8). The type-H consumers’ outside option $\max\left\{v_{H}-p_{0},0\right\}=0$ as $p_{0}=v_{H}$. So the intermediary’s problem is simplified into
$\max_{f,\pi}\text{ }f_{H}\mu_{0}+f_{L}\left(1-\mu_{0}\right)-\mu_{0}\pi\left(l|H\right)\left(v_{H}-v_{L}\right)+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}$
[…]
which are
$f_{H}-f_{L}\leq\left(H-L\right)^{2}-\left[\pi\left(l|L\right)-\pi\left(l|H\right)\right]\left(v_{H}-v_{L}\right)$ ($IC_{H}$)
$-\left(H-L\right)^{2}\leq f_{H}-f_{L}$ ($IC_{L}$)
$f_{H}\leq V+\pi\left(l|H\right)\left(v_{H}-v_{L}\right)$ ($IR_{H}$)
$f_{L}\leq V$ ($IR_{L}$)
$\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)\leq\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)$ ($OB_{L}$)
$\pi\left(l|L\right),\pi\left(l|H\right)\in\left[0,1\right]$ ($FE$)
Note that at least one of $IR_{L}$ (15) and $IR_{H}$ (17) is binding. Otherwise, increasing $f_{L}$ and $f_{H}$ simultaneously by equal size will not affect $IC_{H}$ (12) and $IC_{L}$ (14) and will increase the intermediary’s profit. The following Lemma excludes the case of only $IR_{H}$ (17) being binding.
Lemma 4.
If $IR_{H}$ (17) is binding, then $IR_{L}$ (15) is also binding.
Proof.
Given the binding $IR_{H}$: $f_{H}=V+\pi\left(l|H\right)\Delta v$, then at least one of $IC_{L}$ and $IR_{L}$ is binding. Otherwise, increasing $f_{L}$ increases the objective without affecting constraints $IC_{H},IC_{L},IR_{L}$. Suppose $IC_{L}$ is binding: $f_{H}-f_{L}=-\delta^{2}$. From binding $IR_{H}$ and $IC_{L}$, $f_{L}=V+\pi\left(l|H\right)\Delta v+\delta^{2}>V$ which violates $IR_{L}$. Hence, $IC_{L}$ is not binding and $IR_{L}$ must be binding. ∎
Hence, there are two cases following the above Lemma: $(1)$ only $IR_{L}$ is binding; $(2)$ both $IR_{L}$ and $IR_{H}$ are binding. Now we show Proposition 1 by three steps: Step 1 shows that it’s not optimal that only $IR_{L}$ is binding; Step 2 simplifies the intermediary’s problem with binding $IR_{L}$ and $IR_{H}$; Step 3 solves the problem.
Step 1.
Suppose $IR_{L}$ is binding and $IR_{H}$ is not binding. As $IR_{H}$ is not binding, $IC_{H}$ must be binding. Otherwise, increasing $f_{H}$ will not affect $IR_{H}$, $IC_{H}$, $IC_{L}$, $IR_{L}$ while increases the intermediary’s profit. So $IC_{H}$ must be binding: $f_{H}-V=\delta^{2}-\left[\pi\left(l|L\right)-\pi\left(l|H\right)\right]\Delta v$. From $IC_{L}$
$-\delta^{2}\leq f_{H}-V=\delta^{2}-\left[\pi\left(l|L\right)-\pi\left(l|H\right)\right]\Delta v$
and from (non-binding) $IR_{H}$
$f_{H}-V=\delta^{2}-\left[\pi\left(l|L\right)-\pi\left(l|H\right)\right]\Delta v\leq\pi\left(l|H\right)\Delta v$
so $\pi\left(l|L\right)-\pi\left(l|H\right)\leq\frac{2\delta^{2}}{\Delta v}$ and $\pi\left(l|L\right)\geq\frac{\delta^{2}}{\Delta v}$.
Plugging the binding $IR_{L}$ and $IC_{H}$: $f_{L}=V$ and $f_{H}=V+\delta^{2}-\left[\pi\left(l|L\right)-\pi\left(l|H\right)\right]\Delta v$ into the objective, ignoring the nonbinding $IR_{H}$, the problem is to maximize
$V+\mu_{0}\delta^{2}-\mu_{0}\left[\pi\left(l|L\right)-\pi\left(l|H\right)\right]\Delta v-\mu_{0}\pi\left(l|H\right)\left(v_{H}-v_{L}\right)+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}$
$=$ $V+\mu_{0}\delta^{2}+\pi\left(l|L\right)\left[v_{L}-\mu_{0}v_{H}\right]$
subject to $\pi\left(l|L\right)-\pi\left(l|H\right)\leq\frac{2\delta^{2}}{\Delta v}$. The solution is $\pi\left(l|L\right)=0$ because $v_{L}-\mu_{0}v_{H}<0$. And $\pi\left(l|H\right)=0$ by $OB_{L}$: $\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)\leq\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)$. But the nonbinding $IR_{H}$ requires $\pi\left(l|L\right)\geq\frac{\delta^{2}}{\Delta v}$, which does not hold at $\pi\left(l|L\right)=0$. Hence, it’s not feasible that $IR_{L}$ is binding while $IR_{H}$ is not binding.
Step 2.
Following the first step, it must be that both $IR_{L}$ and $IR_{H}$ are binding: $f_{H}=V+\pi\left(l|H\right)\Delta v$, $f_{L}=V$. Then from $IC_{L}$
$-\delta^{2}\leq\pi\left(l|H\right)\Delta v$
and from $IC_{H}$
$\pi\left(l|L\right)\Delta v-\delta^{2}\leq 0$
so $\pi\left(l|L\right)\leq\frac{\delta^{2}}{\Delta v}$. By plugging the binding $IR_{L}$ and $IR_{H}$ into the objective, the intermediary chooses $\pi\left(l|L\right)$ and $\pi\left(l|H\right)$ to maximize
$V+\mu_{0}\pi\left(l|H\right)\Delta v-\mu_{0}\pi\left(l|H\right)\Delta v+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}$
subject to $OB_{L}$, $\pi\left(l|L\right)\leq\frac{\delta^{2}}{\Delta v}$ and $\pi\left(l|L\right),\pi\left(l|H\right)\in\left[0,1\right]$.
Step 3.
Hence, the solution depends on $\frac{\delta^{2}}{\Delta v}$. When $\frac{\delta^{2}}{\Delta v}\geq 1$, $\pi\left(l|L\right)=1$ and $\pi\left(l|H\right)$ is arbitrary if it satisfies $OB_{L}$: $\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\geq(1-\frac{v_{L}}{v_{H}})\mu_{0}\pi\left(l|H\right)$. The intermediary’s payoff is $V+\left(1-\mu_{0}\right)v_{L}$.
When $\frac{\delta^{2}}{\Delta v}<1$, $\pi\left(l|L\right)=\frac{\delta^{2}}{\Delta v}$ and $\pi\left(l|H\right)$ is arbitrary if it satisfies $OB_{L}$: $\pi(l|H)\leq\frac{C\delta^{2}}{\Delta v}$. The corresponding intermediary payoff is $V+\left(1-\mu_{0}\right)v_{L}\Delta v$.
A.4 Proof of Proposition 2
Proof.
When $\mu_{0}<\frac{v_{L}}{v_{H}}$. The obedience constraint $OB_{L}$ (9) is redundant. This is because $OB_{H}$ (8) and $OB_{L}$ (9) are respectively
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)$$\geq$ $\frac{v_{L}}{v_{H}}-\mu_{0}$
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)$$\geq$ 00 0
and $\frac{v_{L}}{v_{H}}-\mu_{0}>0$. The intermediary’s problem is simplified into
$\max_{f,\pi}\text{ }f_{H}\mu_{0}+f_{L}\left(1-\mu_{0}\right)-\mu_{0}\pi\left(l|H\right)\left(v_{H}-v_{L}\right)+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}+\mu_{0}v_{H}-v_{L}$
s.t. (8) (12) (14) (15) (18) (16)
Note that at least one of $IR_{L}$ and $IR_{H}$ (18) is binding. Otherwise, increasing $f_{L}$ and $f_{H}$ simultaneously by equal size will not affect $IC_{H}$, $IC_{L}$, $IR_{H}$ (18), $IR_{L}$ and will increase the intermediary’s payoff. In addition, the following Lemma excludes the case of only $IR_{L}$ being binding.
Lemma 5.
If $IR_{L}$ (15) is binding, then $IR_{H}$ (18) is also binding.
Proof.
Given $IR_{L}$ is binding $f_{L}=V$, we need to show (18) is binding. Suppose (18) is not binding. Then $IC_{H}$ must be binding. Otherwise, increasing $f_{H}$ will not affect $IC_{H}$, (18), $IC_{L}$ and increase the intermediary’s payoff. So $IC_{H}$ is binding and it gives $f_{H}-V=\delta^{2}-\left[\pi\left(l|L\right)-\pi\left(l|H\right)\right]\Delta v$. From (18)
$f_{H}-V=\delta^{2}-\left[\pi\left(l|L\right)-\pi\left(l|H\right)\right]\Delta v\leq\left[\pi\left(l|H\right)-1\right]\Delta v$
so $\pi\left(l|L\right)\geq 1+\frac{\delta^{2}}{\Delta v}$, which contradicts with $\pi\left(l|L\right)\in[0,1]$. Hence, (18) is binding. ∎
Hence, there are two cases: $(1)$ both $IR_{L}$ and $IR_{H}$ (18) are binding; $(2)$ only $IR_{H}$ (18) is binding. Now we are ready to prove Proposition 2.
Step 1.
In this step, we show that both $IR_{L}$ and $IR_{H}$ (18) are binding. Suppose not, then only $IR_{H}$ (18) is binding: $f_{H}=V+\left[\pi\left(l|H\right)-1\right]\Delta v$. As $IR_{L}$ is not binding, $IC_{L}$ must be binding. Otherwise, increasing $f_{L}$ will not affect $IR_{L}$, $IC_{L}$, $IC_{H}$ and will increase the intermediary’s payoff. So $IC_{L}$ is binding. The binding $IC_{L}$ and $IR_{H}$ (18) give $f_{L}=f_{H}+\delta^{2}=V+\left[\pi\left(l|H\right)-1\right]\Delta v+\delta^{2}$. From $IC_{H}$
$-\delta^{2}\leq\delta^{2}-\left[\pi\left(l|L\right)-\pi\left(l|H\right)\right]\Delta v$
and from (non-binding) $IR_{L}$
$V+\left[\pi\left(l|H\right)-1\right]\Delta v+\delta^{2}<V$
or we rewrite $IC_{H}$ and $IR_{L}$ as
$\pi\left(l|L\right)-\pi\left(l|H\right)$$\leq$ $\frac{2\delta^{2}}{\Delta v}$
$\pi\left(l|H\right)$$\leq$ $1-\frac{\delta^{2}}{\Delta v}$
Plug $f_{L}$ and $f_{H}$ from the binding $IR_{H}$ and $IC_{L}$ into the objective and ignore the nonbinding constraint, the intermediary maximizes
$\mu_{0}\left[V+\left(\pi\left(l|H\right)-1\right)\Delta v\right]+(1-\mu_{0})\left[V+\left(\pi\left(l|H\right)-1\right)\Delta v+\delta^{2}\right]$
$+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}-\mu_{0}\pi\left(l|H\right)\Delta v+\mu_{0}v_{H}-v_{L}$
$=$ $V+\left[\pi\left(l|H\right)-1\right]\Delta v+(1-\mu_{0})\delta^{2}+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}-\mu_{0}\pi\left(l|H\right)\Delta v+\mu_{0}v_{H}-v_{L}$
$=$ $V+(1-\mu_{0})\delta^{2}-\left(1-\mu_{0}\right)v_{H}+\left(1-\mu_{0}\right)\Delta v\pi\left(l|H\right)+\left(1-\mu_{0}\right)v_{L}\pi\left(l|L\right)$
subject to $OB_{H}$ and $\pi\left(l|L\right)-\pi\left(l|H\right)\leq\frac{2\delta^{2}}{\Delta v}$.
The solution to this reduced problem is $\pi\left(l|L\right)=1$ and $\pi\left(l|H\right)=1$. Check the nonbinding $IR_{L}$: $\pi\left(l|H\right)=1<1-\frac{\delta^{2}}{\Delta v}$, which does not hold. So it is not optimal that only $IR_{H}$ (18) is binding. Hence, $IR_{L}$ and $IR_{H}$ (18) are binding.
Step 2.
Following the first step, it must be both $IR_{L}$ and $IR_{H}$ (18) are binding: $f_{H}=V+\left[\pi\left(l|H\right)-1\right]\Delta v$, $f_{L}=V$. Then plugging $f_{L}$, $f_{H}$ into $IC_{H}$
$\left[\pi\left(l|H\right)-1\right]\Delta v\leq\delta^{2}-\left[\pi\left(l|L\right)-\pi\left(l|H\right)\right]\Delta v$
and into $IC_{L}$
$\left[1-\pi\left(l|H\right)\right]\Delta v-\delta^{2}\leq 0$
or
$\pi\left(l|L\right)$$\leq$ $1+\frac{\delta^{2}}{\Delta v}$
$\pi\left(l|H\right)$$\geq$ $1-\frac{\delta^{2}}{\Delta v}$
By plugging $f_{H}=V+\left[\pi\left(l|H\right)-1\right]\Delta v$, $f_{L}=V$ into the objective, the intermediary chooses $\pi\left(l|L\right)$ and $\pi\left(l|H\right)$ to maximize
$V+\mu_{0}\left[\pi\left(l|H\right)-1\right]\Delta v-\mu_{0}\pi\left(l|H\right)\Delta v+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}+\mu_{0}v_{H}-v_{L}$
$=$ $V+\left(1-\mu_{0}\right)v_{L}\pi\left(l|L\right)-\left(1-\mu_{0}\right)v_{L}$
subject to $OB_{H}$ and
$\pi\left(l|H\right)$$\geq$ $1-\frac{\delta^{2}}{\Delta v}$
$\pi\left(l|L\right)$$\leq$ $1+\frac{\delta^{2}}{\Delta v}$
Step 3.
The solution to the reduced problem is $\pi\left(l|L\right)=1$, and $\pi\left(l|H\right)$ depends on $\Delta v$. When $\frac{\delta^{2}}{\Delta v}\geq 1$, $\pi\left(l|H\right)\in\left[0,1\right]$. The profit from the service is $V-\mu_{0}\Delta v\left(1-\pi\left(l|H\right)\right)$ and the profit from the data trade is $\mu_{0}\Delta v\left(1-\pi\left(l|H\right)\right)$. So its total payoff is $V$.
When $\frac{\delta^{2}}{\Delta v}<1$, $\pi\left(l|H\right)\in\big{[}1-\frac{\delta^{2}}{\Delta v},1\big{]}$ and the payoff is $V$.
A.5 Proof of Corollary 3
Proof.
Suppose $\mu_{0}>\frac{v_{L}}{v_{H}}$ (the other case is similar). To see the full information data trade is optimal if and only if $\frac{\delta^{2}}{\Delta v}\geq 1$. Suppose the data trade is fully revealing $\pi(l|L)=1-\pi(l|H)=1$, $OB_{L}$ is satisfied in this situation. The maximum revenue that the intermediary can achieve under the full revelation data trade is the value of the following problem
$\max_{f_{L},f_{H}}$ $\quad f_{H}\mu_{0}+f_{L}\left(1-\mu_{0}\right)+\left(1-\mu_{0}\right)v_{L}$
s.t. $f_{H}-f_{L}\leq\delta^{2}-\Delta v$
$f_{H}-f_{L}\geq-\delta^{2}$
$f_{H}\leq V$
$f_{L}\leq V$
where the constraints on service fees are from (12) (13) (14) (15). To make the choice set nonempty, it requires $\delta^{2}-\Delta v\geq-\delta^{2}$; otherwise, there is no solution to the problem, i.e., the full revelation data trade is not incentive compatible. The solution of $f_{H}$, $f_{L}$ and the maximum revenue are
  1. (i)
    for $\delta^{2}\geq\Delta v$, then $f_{H}=f_{L}=V$ and the revenue is $V+\left(1-\mu_{0}\right)v_{L}$. The full reveling achieves the maximum payoff for $\frac{\delta^{2}}{\Delta v}\geq 1$.
  2. (ii)
    for $\delta^{2}<\Delta v$, then $f_{L}=V$ and $f_{H}=V+\delta^{2}-\Delta v$ and the revenue is $V+\left(1-\mu_{0}\right)v_{L}+\mu_{0}\left(\delta^{2}-\Delta v\right)$. The full reveling is not optimal because the intermediary’s total revenue from full revelation is lower than that from the equilibrium one
    $V+\left(1-\mu_{0}\right)v_{L}+\mu_{0}\left(\delta^{2}-\Delta v\right)-[V+\left(1-\mu_{0}\right)v_{L}\frac{\delta^{2}}{\Delta v}]$
    $=$ $(\delta^{2}-\Delta v)\frac{\mu_{0}v_{H}-v_{L}}{\Delta v}<0$
    when $\delta^{2}<\Delta v$ and $\mu_{0}>\frac{v_{L}}{v_{H}}$.
A.6 Proof of Proposition 3
Proof.
Without a data market, the intermediary maximizes its service revenue by choosing the service provisions and fees
$\max_{f,x}\text{ }$ $f_{H}\mu_{0}+f_{L}\left(1-\mu_{0}\right)$
s.t. $V-\left(x_{H}-H\right)^{2}-f_{H}\geq V-\left(x_{L}-H\right)^{2}-f_{L}$
$V-\left(x_{L}-L\right)^{2}-f_{L}\geq V-\left(x_{H}-L\right)^{2}-f_{H}$
$V-\left(x_{H}-H\right)^{2}-f_{H}\geq 0$
$V-\left(x_{L}-L\right)^{2}-f_{L}\geq 0$
With the assumption $\mathcal{X}=\left\{L,H\right\}$, $x_{L}=L$ and $x_{H}=H$ by Lemma 3. Then the constraints turn into
$-f_{H}$ $\geq-\delta^{2}-f_{L}$
$-f_{L}$ $\geq-\delta^{2}-f_{H}$
$f_{H}$ $\leq V$
$f_{L}$ $\leq V$
Hence, $f_{L}=f_{H}=V$.
A.7 Proof of Proposition 5
Proof.
When $\mu_{0}>\frac{v_{L}}{v_{H}}$. From $OB_{H}$ (8) and $OB_{L}$ (9)
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)$$\geq$ $\frac{v_{L}}{v_{H}}-\mu_{0}$
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)$$\geq$ 00 0
The obedience constraint for the producer to charge a high price $OB_{H}$ (8) is redundant, we can ignore (8). The type-H consumers’ outside option $\max\left\{v_{H}-p_{0},0\right\}=0$. So the intermediary’s problem is simplified into
$\max_{f,\pi}\text{ }f_{H}\mu_{0}+f_{L}\left(1-\mu_{0}\right)-\mu_{0}\pi\left(l|H\right)\left(v_{H}-v_{L}\right)+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}$
[…]
which are
$V-\left(x_{H}-H\right)^{2}-f_{H}+\pi\left(l|H\right)\Delta v\geq V-\left(x_{L}-H\right)^{2}-f_{L}+\pi\left(l|L\right)\Delta v$ ($IC_{H}$ )
$V-\left(x_{L}-L\right)^{2}-f_{L}\geq V-\left(x_{H}-L\right)^{2}-f_{H}$ ($IC_{L}$)
$V-\left(x_{H}-H\right)^{2}-f_{H}+\pi\left(l|H\right)\Delta v\geq 0$ ($IR_{H}$)
$V-\left(x_{L}-L\right)^{2}-f_{L}\geq 0$ ($IR_{L}$)
$\pi\left(l|H\right)\leq C\pi\left(l|L\right)$ ($OB_{L}$)
$\pi\left(l|L\right),\pi\left(l|H\right)\in\left[0,1\right]$ ($FE$)
Note that at least one of $IR_{L}$ (15) and $IR_{H}$ (17) is binding. Otherwise, both of $IR_{L}$ (15) and $IR_{H}$ (17) are not binding. So increasing $f_{L}$ and $f_{H}$ simultaneously by equal size will not affect $IC_{H}$ (12) and $IC_{L}$ (14), and will increase the intermediary’s profit. The following Lemma excludes the case of only $IR_{H}$ (17) being binding.
Lemma 6.
If $IR_{H}$ (17) is binding, then $IR_{L}$ (15) is also binding.
Proof.
Given the binding $IR_{H}$: $V-(x_{H}-H)^{2}-f_{H}+\pi\left(l|H\right)\Delta v=0$, then at least one of $IC_{L}$ and $IR_{L}$ is binding. Otherwise, increasing $f_{L}$ increases the objective without affecting constraints $IC_{H}$, $IC_{L}$, $IR_{L}$.
Suppose $IC_{L}$ is binding: $V-(x_{L}-L)^{2}-f_{L}=V-(x_{H}-L)^{2}-f_{H}$. From the binding $IR_{H}$ and $IC_{L}$,
$V-(x_{L}-L)^{2}-f_{L}$
$=$ $V-(x_{H}-L)^{2}-V+(x_{H}-H)^{2}-\pi\left(l|H\right)\Delta v$
$=$ $(L-H)(2x_{H}-L-H)-\pi\left(l|H\right)\Delta v$
$<$ 00 0
which violates $IR_{L}$. Hence, $IC_{L}$ cannot be binding, and $IR_{L}$ must be binding. ∎
Hence, there are two cases following the above Lemma: $(1)$ $IR_{L}$ is binding, $IR_{H}$ is not binding; $(2)$ both $IR_{L}$ and $IR_{H}$ are binding. Now we show Proposition 1 by three steps: Step 1 shows that it’s not optimal that only $IR_{L}$ is binding by contradiction; Step 2 simplifies the intermediary’s problem with binding $IR_{L}$ and $IR_{H}$; Step 3 solves the problem.
Step 1.
Suppose $IR_{L}$ is binding, $IR_{H}$ is not binding. As $IR_{H}$ is not binding, $IC_{H}$ must be binding. Otherwise, increasing $f_{H}$ will not affect $IR_{H}$, $IC_{H}$, $IC_{L}$, $IR_{L}$ while increases the intermediary’s profit. So $IC_{H}$ is binding:
$V-\left(x_{H}-H\right)^{2}-f_{H}+\pi\left(l|H\right)\Delta v=V-\left(x_{L}-H\right)^{2}-f_{L}+\pi\left(l|L\right)\Delta v$ (19)
also recall the binding $IR_{L}$: $V-(x_{L}-L)^{2}-f_{L}=0$, we have
$f_{L}=V-(x_{L}-L)^{2}$ (20)
Plugging (20) into (19) to get
$f_{H}=V+\left(x_{L}-H\right)^{2}-(x_{L}-L)^{2}-\left(x_{H}-H\right)^{2}-\pi\left(l|L\right)\Delta v+\pi\left(l|H\right)\Delta v$ (21)
Using $f_{L}$ in (20) and $f_{H}$ in (21), $IC_{L}$ becomes
00 0 $\geq$ $V-(x_{H}-L)^{2}-(x_{L}-H)^{2}+(x_{H}-H)^{2}+(x_{L}-L)^{2}-\left[\pi\left(l|L\right)+\pi\left(l|H\right)\right]\Delta v$
$=$ $(x_{L}-L)^{2}-(x_{L}-H)^{2}+(x_{H}-H)^{2}-(x_{H}-L)^{2}-\left[\pi\left(l|L\right)+\pi\left(l|H\right)\right]\Delta v$
and (non-binding) $IR_{H}$ is
$(x_{L}-L)^{2}-(x_{L}-H)^{2}+\pi(l|L)\Delta v\geq 0$ (22)
Plugging (20) and (21) into the objective function, the problem is to maximize
$\quad\mu_{0}\left[(x_{L}-H)^{2}-(x_{H}-H)^{2}+(\pi(l|H)-\pi(l|L))\Delta v+V-(x_{L}-L)^{2}\right]$ (23)
$+(1-\mu_{0})\left[V-(x_{L}-L)^{2}\right]-\mu_{0}\pi(l|H)\Delta v+(1-\mu_{0})\pi(l|L)v_{L}$
$=$ $V+\mu_{0}\left[(x_{L}-H)^{2}-(x_{H}-H)^{2}\right]-(x_{L}-L)^{2}+\pi(l|L)(v_{L}-\mu_{0}v_{H})$
subject to $IC_{L}$ and $OB_{L}$.
We note that $IC_{L}$ is binding. To see this by contradiction, suppose the constraints ($IC_{L}$) are not binding. We have $\pi(l|L)=0$ as $v_{L}-\mu_{0}v_{H}<0$. Then $\pi(l|H)=0$ and $x_{L}=\frac{L-\mu_{0}H}{1-\mu_{0}}$ and $x_{H}=H$. But this contradicts with $IR_{H}$ in (22) as
$\left(\frac{L-\mu_{0}H}{1-\mu_{0}}-L\right)^{2}-\left(\frac{L-\mu_{0}H}{1-\mu_{0}}-H\right)^{2}<0.$
Using the binding $IC_{L}$, i.e.,
$(x_{L}-L)^{2}-(x_{L}-H)^{2}+(x_{H}-H)^{2}-(x_{H}-L)^{2}-\left[\pi\left(l|L\right)+\pi\left(l|H\right)\right]\Delta v=0$
we have
$2\delta(x_{L}-x_{H})=\left[\pi(l|H)-\pi(l|L)\right]\Delta v$
from which by substituting out $x_{H}$, the objective becomes
$V+\mu_{0}\left[(x_{L}-H)^{2}-(x_{H}-H)^{2}\right]-(x_{L}-L)^{2}+\pi(l|L)(v_{L}-\mu_{0}v_{H})$
$=$ $V+\mu_{0}\left[(x_{L}-H)^{2}-\left(x_{L}+\frac{(\pi(l|L)-\pi(l|H))\Delta v}{2\delta}-H\right)^{2}\right]-(x_{L}-L)^{2}+\pi(l|L)(v_{L}-\mu_{0}v_{H})$
We have two cases depending on whether $OB_{L}$ is binding or not and the following shows that neither case can be optimal.
(1) $OB_{L}$ is not binding. Then the first order conditions with $x_{L}$, $\pi(l|H)$, $\pi(l|L)$ are
$x_{L}=L-\frac{\mu_{0}\Delta v\left(\pi(l|L)-\pi(l|H)\right)}{2\delta}$
$-2\mu_{0}\left(x_{H}-H\right)\frac{-\Delta v}{2\delta}\geq 0$
$-2\mu_{0}\left(x_{H}-H\right)\frac{\Delta v}{2\delta}+(v_{L}-\mu_{0}v_{H})\leq 0$
Suppose $x_{H}>H$. Then $\pi(l|H)=1$ and $\pi(l|L)=0$, which induces a contradiction with $OB_{L}$, $\pi(l|H)\leq\pi(l|L)$. Hence, $x_{H}=H$. Then $\pi(l|L)=0$. From the binding $IC_{L}$, $02\delta(x_{L}-x_{H})=(\pi(l|H)-\pi(l|L))\Delta v<$ 0, which implies $\pi(l|H)<0$, a contradiction.
(2) $OB_{L}$ is binding: $\pi(l|H)=C\pi(l|L)$. Then the objective function becomes
$V+\mu_{0}\left[(x_{L}-H)^{2}-(x_{L}+\frac{(\pi(l|L)-\pi(l|H))\Delta v}{2\delta}-H)^{2}\right]-(x_{L}-L)^{2}+\pi(l|L)(v_{L}-\mu_{0}v_{H})$
$=$ $V+\mu_{0}\left[(x_{L}-H)^{2}-(x_{L}+\frac{(\pi(l|L)-C\pi(l|L))\Delta v}{2\delta}-H)^{2}\right]-(x_{L}-L)^{2}+\pi(l|L)(v_{L}-\mu_{0}v_{H})$
Its first order derivative with $\pi(l|L)$ is $-2\mu_{0}\left(x_{H}-H\right)\frac{(1-C)\Delta v}{2\delta}+(v_{L}-\mu_{0}v_{H})<0$, so $\pi(l|L)=0$ and then $\pi(l|H)=0$ by the binding $OB_{L}$. Hence, $x_{L}=x_{H}=L$ following the first order condition for $x_{L}$, i.e., $\mu_{0}(x_{L}-x_{H})-x_{L}+L=0$ and the binding $IC_{L}$: $02\delta(x_{L}-x_{H})=$ 0. It is a contradiction.
In sum, Step 1 shows by contradiction that it is not feasible that $IR_{L}$ is binding while $IR_{H}$ is not binding.
Step 2.
Following the first step, it must be that both $IR_{L}$ and $IR_{H}$ are binding:
$f_{H}$ $=$ $V-\left(x_{H}-H\right)^{2}+\pi\left(l|H\right)\Delta v$
$f_{L}$ $=$ $V-\left(x_{L}-L\right)^{2}$
By plugging $f_{L}$ and $f_{H}$ into the objective, the intermediary chooses $x_{H}$, $x_{L}$ and $\pi\left(l|L\right)$, $\pi\left(l|H\right)$ to maximize
$V-\mu_{0}\left(x_{H}-H\right)^{2}-\left(1-\mu_{0}\right)\left(x_{L}-L\right)^{2}+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}$
subject to $OB_{L}$, $IC_{L}$, $IC_{H}$ and $\pi\left(l|L\right),\pi\left(l|H\right)\in\left[0,1\right]$.
By plugging $f_{L}$ and $f_{H}$ into $IC_{L}$ and $IC_{H}$, we have the following $OB_{L}$, $IC_{L}$, $IC_{H}$
$\left(x_{H}-H\right)^{2}-\left(x_{H}-L\right)^{2}$ $\leq\pi\left(l|H\right)\Delta v$ ($IC_{L}$)
$\left(x_{L}-L\right)^{2}-\left(x_{L}-H\right)^{2}$ $\leq-\pi\left(l|L\right)\Delta v$ ($IC_{H}$ )
$\pi\left(l|H\right)$ $\leq C\pi\left(l|L\right)$ ($OB_{L}$)
Step 3.
Solving the simplified problem. Note that $x_{H}=H$, and $IC_{L}$ and $OB_{L}$ are redundant. If $IC_{H}$ is not binding, then $x_{L}=L$, $\pi(l|L)=1$. $IC_{H}$ requires $\Delta v\leq\delta^{2}$.
So if $\Delta v>\delta^{2}$, $IC_{H}$ is binding, i.e., $\left(x_{L}-L\right)^{2}-\left(x_{L}-H\right)^{2}=-\pi\left(l|L\right)\Delta v$ from which
$(2x_{L}-L-H)\delta=-\pi(l|L)\Delta v$
or
$x_{L}=\frac{L+H}{2}-\frac{\pi(l|L)\Delta v}{2\delta}$
Substituting out $x_{L}$ in the objective function, the intermediary chooses $\pi\left(l|L\right)\in\left[0,1\right]$ to maximize
$-\left[\frac{\delta}{2}-\frac{\pi(l|L)\Delta v}{2\delta}\right]^{2}+\pi\left(l|L\right)v_{L}$
The solution to this problem is
  • if $\frac{\delta^{2}}{\Delta v}\leq\frac{\Delta v}{v_{L}+v_{H}}$, then $\pi(l|L)=\frac{\delta^{2}}{\Delta v}\frac{v_{H}+v_{L}}{\Delta v}$ and $x_{L}=L-\frac{\delta}{\Delta v}v_{L}$
  • if $\frac{\Delta v}{v_{L}+v_{H}}<\frac{\delta^{2}}{\Delta v}<1$, then $\pi(l|L)=1$ and $x_{L}=L-\frac{\Delta v-\delta^{2}}{2\delta}$
Summarizing the above results,
  • if $\frac{\delta^{2}}{\Delta v}\geq 1$, then $\pi(l|L)=1$ and $x_{L}=L$
  • if $\frac{\delta^{2}}{\Delta v}\leq\frac{\Delta v}{v_{L}+v_{H}}$, then $\pi(l|L)=\frac{\delta^{2}}{\Delta v}\frac{v_{H}+v_{L}}{\Delta v}$ and $x_{L}=L-\frac{\delta}{\Delta v}v_{L}$
  • if $\frac{\Delta v}{v_{L}+v_{H}}<\frac{\delta^{2}}{\Delta v}<1$, then $\pi(l|L)=1$ and $x_{L}=L-\frac{\Delta v-\delta^{2}}{2\delta}$
The intermediary’s total revenue is
$\begin{cases}V+(1-\mu_{0})v_{L}&\mbox{ if }\frac{\delta^{2}}{\Delta v}\geq 1\\ V-(1-\mu_{0})\left(\frac{\Delta v-\delta^{2}}{2\delta}\right)^{2}+(1-\mu_{0})v_{L}&\mbox{ if }\frac{\Delta v}{v_{L}+v_{H}}\leq\frac{\delta^{2}}{\Delta v}<1\\ V+(1-\mu_{0})\left(\frac{\delta}{\Delta v}\right)^{2}v_{L}v_{H}&\mbox{ if }\frac{\delta^{2}}{\Delta v}<\frac{\Delta v}{v_{L}+v_{H}}\end{cases}$
which is continuous in $\delta^{2}/\Delta v$ and hence is consistent with Berge’s maximum theorem.
A.8 Additional Results on General Service
Proposition 7.
Suppose the service space $\mathcal{X}=\mathbb{R}$ and $\mu_{0}<\frac{v_{L}}{v_{H}}$. In the optimal mechanism, there is no distortion in service provisions $x_{L}=L$, $x_{H}=H$, the service fees are $f_{L}=V$, $f_{H}=V+[\pi(l|H)-1]\Delta v$, and the intermediary’s data trade is
$\pi\left(l|L\right)=1,\text{ }\pi\left(l|H\right)\in\Big{[}\max\Big{\{}0,1-\frac{\delta^{2}}{\Delta v}\Big{\}},1\Big{]}$
and the data trade fee is
$T=\mu_{0}[1-\pi\left(l|H\right)]\Delta v.$
In this case, social welfare is
$V+\left(1-\mu_{0}\right)v_{L}+\mu_{0}v_{H}.$
This actually goes back to the binary analysis in Proposition 2. The following proves Proposition 7.
Proof.
When $\mu_{0}<\frac{v_{L}}{v_{H}}$, the obedience constraint $OB_{L}$ is redundant, and the type-H consumers’ outside option max $\left\{v_{H}-p_{0},0\right\}=v_{H}-v_{L}$. Hence, the intermediary’s problem is simplified into
$\max_{f,x,\pi,T}f_{H}\mu_{0}+f_{L}\left(1-\mu_{0}\right)+T$
s.t.
$V-\left(x_{H}-H\right)^{2}-f_{H}+\pi\left(l|H\right)\left(v_{H}-v_{L}\right)$
$\hskip 56.9055pt\geq V-\left(x_{L}-H\right)^{2}-f_{L}+\pi\left(l|L\right)\left(v_{H}-v_{L}\right)$ ($IC_{H}$ )
$V-\left(x_{L}-L\right)^{2}-f_{L}\geq V-\left(x_{H}-L\right)^{2}-f_{H}$ ($IC_{L}$)
$V-\left(x_{H}-H\right)^{2}-f_{H}+\pi\left(l|H\right)\left(v_{H}-v_{L}\right)\geq v_{H}-v_{L}$ ($IR_{H}$)
$V-\left(x_{L}-L\right)^{2}-f_{L}\geq 0$ ($IR_{L}$)
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)\geq\frac{v_{L}}{v_{H}}-\mu_{0}$ ($OB_{H}$)
$R\left(\pi\right)-R_{0}\geq T$ ($IR_{S}$)
$\pi\left(l|L\right),\pi\left(l|H\right)\in\left[0,1\right]$ ($FE$)
Lemma 7.
If $IR_{L}$ (15) is binding, then $IR_{H}$ (18) is also binding.
Proof.
Given $IR_{L}$ is binding $f_{L}=V-\left(x_{L}-L\right)^{2}$. First, at least one of $IC_{H}$ and $IR_{H}$ is binding. Suppose not, both are not binding. Increasing $f_{H}$ will not affect $IC_{H}$, $IR_{H}$ (18), $IC_{L}$ and increase the intermediary’s payoff.
Second, $IC_{H}$ is not binding. To see this, suppose $IC_{H}$ is binding, then it gives
$V-\left(x_{H}-H\right)^{2}-f_{H}+\pi\left(l|H\right)\left(v_{H}-v_{L}\right)=V-\left(x_{L}-H\right)^{2}-f_{L}+\pi\left(l|L\right)\left(v_{H}-v_{L}\right)$
Together with $f_{L}=V-\left(x_{L}-L\right)^{2}$, we have
$f_{H}$ $=V+\left(x_{L}-H\right)^{2}-\left(x_{H}-H\right)^{2}-\left(x_{L}-L\right)^{2}-(\pi\left(l|L\right)-\pi\left(l|H\right))\Delta v$
$\leq V-\left(x_{H}-H\right)^{2}+\pi\left(l|H\right)\Delta v-\Delta v$
where the inequality follows from $IR_{H}$. So
$2x_{L}\geq L+H$
leading a contradiction with $x_{L}\leq L.$ Hence, $IC_{H}$ is not binding.
Thus, by the above two points, $IR_{H}$ is binding. ∎
In addition, we note that at least one of $IR_{L}$ and $IR_{H}$ is binding. Otherwise, increasing $f_{L}$ and $f_{H}$ simultaneously by equal size will not affect $IC_{H}$, $IC_{L}$, $IR_{H}$, $IR_{L}$ and will increase the intermediary’s payoff.
Hence, by the above lemma, there are two cases: $(1)$ both $IR_{L}$ and $IR_{H}$ are binding; $(2)$ only $IR_{H}$ is binding. The following Step 1 shows that Case (2) is not feasible by contradiction, and Step 2 solves the optimal mechanism under Case (1).
Step 1.
We show that only $IR_{H}$ (18) is binding is not feasible. Suppose not, only (18) is binding: $f_{H}=V-\left(x_{H}-H\right)^{2}+\left[\pi\left(l|H\right)-1\right]\Delta v$. As $IR_{L}$ is not binding, $IC_{L}$ must be binding. Otherwise, increasing $f_{L}$ will not affect $IR_{L}$, $IC_{L}$, $IC_{H}$ and will increase intermediary’s payoff. So the binding (18) and $IC_{L}$ give
$f_{H}$ $=$ $V-\left(x_{H}-H\right)^{2}+(\pi\left(l|H\right)-1)\Delta v$
$f_{L}$ $=$ $V-\left(x_{L}-L\right)^{2}+\left(x_{H}-L\right)^{2}-\left(x_{H}-H\right)^{2}+\left(\pi\left(l|H\right)-1\right)\Delta v$
Then $IC_{H}$ is
$2\left(x_{L}-x_{H}\right)\leq\pi\left(l|H\right)-\pi\left(l|L\right)$
and (non-binding) $IR_{L}$ is
$2x_{H}-v_{L}-v_{H}<1-\pi\left(l|H\right)$
The intermediary’s problem is
$\max_{x,\pi}V-\left(x_{H}-{H}\right)^{2}+\pi$ $(l|H)\left(1-\mu_{0}\right)\Delta v+\left[\left(x_{H}-{L}\right)^{2}-\left(x_{L}-{L}\right)^{2}\right]\left(1-\mu_{0}\right)$
$+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}-\left(1-\mu_{0}\right)v_{H}$
subject to $OB_{H}$ and $IC_{H}$
$2\left(x_{L}-x_{H}\right)\leq\pi\left(l|H\right)-\pi\left(l|L\right)$
and $\pi\left(l|L\right)$, $\pi\left(l|H\right)\in\left[0,1\right]$. If $IC_{H}$ is not binding and ignore the constraint, the solution is
$x_{L}=L,x_{H}=\frac{H-\left(1-\mu_{0}\right)L}{\mu_{0}},\pi\left(l|L\right)=\pi\left(l|H\right)=1$
which satisfies $IC_{H}$ but doesn’t satisfy $IR_{L}$. Hence, $IC_{H}$ is binding:
$2\left(x_{L}-x_{H}\right)=\pi\left(l|H\right)-\pi\left(l|L\right).$ (24)
Plugging (24) into the intermediary’s objective function which maximizes
$V-\left(x_{H}-H\right)^{2}+\pi\left(l|H\right)\left(1-\mu_{0}\right)\Delta v$
$+\Big{[}\left(x_{H}-L\right)^{2}-\Big{(}\frac{\pi\left(l|H\right)-\pi\left(l|L\right)}{2}+x_{H}-L\Big{)}^{2}\Big{]}\left(1-\mu_{0}\right)+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}-\left(1-\mu_{0}\right)v_{H}$
subject to $OB_{H}$, $\pi\left(l|L\right)\in\left[0,1\right]$, $\pi\left(l|H\right)\in\left[0,1\right]$. The first order derivative with respect to $\pi(l|H)$ is positive, $\pi(l|H)=1$.
Suppose $\pi(l|L)\in[0,1]$ is not binding. The first order condition with respect to $\pi(l|L)$ and $x_{H}$ are respectively
$x_{H}-L$ $=$ $-v_{L}-\frac{\pi(l|H)-\pi(l|L)}{2}$
$x_{H}-H$ $=$ $-\frac{\pi\left(l|H\right)-\pi\left(l|L\right)}{2}\left(1-\mu_{0}\right)$
from which
$x_{H}$ $=$ $\frac{H-(L-v_{L})(1-\mu_{0})}{\mu_{0}}$
$\pi(l|L)$ $=$ $1\displaystyle 1+2\times\frac{H-L+v_{L}}{\mu_{0}}>$ 1
So $\pi(l|L)=1$, and then $x_{H}=H$. By (24) $x_{L}=x_{H}=H$ which is a contradiction with $x_{L}\leq L$. Hence, it’s not feasible that only $IR_{H}$ is binding.
Step 2.
Following Step 1, it must be both $IR_{L}$ and $IR_{H}$ are binding:
$f_{H}$ $=$ $V-\left(x_{H}-H\right)^{2}+\pi\left(l|H\right)\Delta v-\Delta v$
$f_{L}$ $=$ $V-\left(x_{L}-L\right)^{2}$
Then from $IC_{H}$
$\Delta v\geq-\left(x_{L}-H\right)^{2}+\left(x_{L}-L\right)^{2}+\pi\left(l|L\right)\Delta v$
and from $IC_{L}$
$0\geq-\left(x_{H}-L\right)^{2}+\left(x_{H}-H\right)^{2}-\pi\left(l|H\right)\Delta v+\Delta v$
equivalently, they are
$1-\pi\left(l|L\right)$$\geq$ $\left(2x_{L}-L-H\right)\frac{\delta}{\Delta v}$
$1-\pi\left(l|H\right)$$\leq$ $\left(2x_{H}-L-H\right)\frac{\delta}{\Delta v}$
By plugging $f_{H}$ and $f_{L}$ into the objective, the intermediary chooses $\pi\left(l|L\right)$ and $\pi\left(l|H\right)$ to maximize
$V-\mu_{0}\left(x_{H}-H\right)^{2}-\left(1-\mu_{0}\right)\left(x_{L}-L\right)^{2}+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}-v_{L}\left(1-\mu_{0}\right)$
subject to $OB_{H}$ and $\pi\left(l|L\right)$, $\pi\left(l|H\right)\in\left[0,1\right]$ and
$1-\pi\left(l|L\right)$$\geq$ $\left(2x_{L}-L-H\right)\frac{\delta}{\Delta v}$
$1-\pi\left(l|H\right)$$\leq$ $\left(2x_{H}-L-H\right)\frac{\delta}{\Delta v}$
Hence, $x_{L}=L$, $x_{H}=H$ and $\pi\left(l|L\right)=1$ and $\pi\left(l|H\right)\in\left[\max\left\{0,1-\frac{\delta^{2}}{\Delta v}\right\},1\right]$.
In sum, when $\mu_{0}<\frac{v_{L}}{v_{H}}$, $x_{L}=L$, $x_{H}=H$, $f_{L}=V$, $f_{H}=V-\left[1-\pi(l|H)\right]\Delta v$, $\pi\left(l|L\right)=1$, $\pi\left(l|H\right)\in\left[\max\left\{0,1-\frac{\delta^{2}}{\Delta v}\right\},1\right]$.
A.9 Proof of Proposition 6
Proof.
For $\mu_{0}<\frac{v_{L}}{v_{H}}$, social welfare without regulation is $V+(1-\mu_{0})v_{L}+\mu_{0}v_{H}$, and social welfare with data banning is $V+v_{L}+\mu_{0}\Delta v$. So banning data trade has no effect on social welfare for $\mu_{0}<\frac{v_{L}}{v_{H}}$.
For $\mu_{0}>\frac{v_{L}}{v_{H}}$, social welfare with data banning is $V+\mu_{0}v_{H}$, and social welfare without regulation
  1. (i)
    if $\frac{\delta^{2}}{\Delta v}\geq 1$: $V+(1-\mu_{0})v_{L}+\mu_{0}v_{H}$
  2. (ii)
    if $\frac{\Delta v}{v_{L}+v_{H}}<\frac{\delta^{2}}{\Delta v}<1$: $V-(1-\mu_{0})\left(\frac{\Delta v-\delta^{2}}{2\delta}\right)^{2}+(1-\mu_{0})v_{L}+\mu_{0}v_{H}$
  3. (iii)
    if $\frac{\delta^{2}}{\Delta v}<\frac{\Delta v}{v_{L}+v_{H}}$: $V+(1-\mu_{0})\left(\frac{\delta}{\Delta v}\right)^{2}v_{L}v_{H}+\mu_{0}v_{H}$
each case is strictly greater than $V+\mu_{0}v_{H}$. This is obvious for the first and third cases. In the second case, $v_{L}-\left(\frac{\Delta v-\delta^{2}}{2\delta}\right)^{2}=\frac{-(\Delta v)^{2}-\delta^{4}+2\delta^{2}(v_{L}+v_{H})}{4\delta^{2}}>0$ when $\frac{\Delta v}{v_{L}+v_{H}}<\frac{\delta^{2}}{\Delta v}<1$. Hence, banning data trade decreases social welfare for $\mu_{0}>\frac{v_{L}}{v_{H}}$.